### v1

# Debiased Sample Selection for Combating Noisy Labels

Abstract
Learning with noisy labels aims to ensure model generalization given a label-corrupted training set. The sample selection strategy achieves promising performance by selecting a label-reliable subset for model training. In this paper, we empirically reveal that existing sample selection methods suffer from both data and training bias that are represented as imbalanced selected sets and accumulation errors in practice, respectively. However, only the training bias was handled in previous studies. To address this limitation, we propose a noIse-Tolerant Expert Model (ITEM) for debiased learning in sample selection. Specifically, to mitigate the training bias, we design a robust network architecture that integrates with multiple experts. Compared with the prevailing double-branch network, our network exhibits better performance of selection and prediction by ensembling these experts while training with fewer parameters. Meanwhile, to mitigate the data bias, we propose a mixed sampling strategy based on two weight-based data samplers. By training on the mixture of two class-discriminative mini-batches, the model mitigates the effect of the imbalanced training set while avoiding sparse representations that are easily caused by sampling strategies. Extensive experiments and analyses demonstrate the effectiveness of ITEM. Our code is available at this url ITEM.
1 Introduction
The remarkable generalization capability of deep neural networks (DNNs) is achieved through training on a large-scale dataset [1, 2]. However, existing training sets are usually collected by online queries [3], crowdsourcing [4], and manual annotations, which could inevitably incur wrong (or noisy) labels [5]. Since DNNs exhibit vulnerability to such low-quality datasets [6], training on the label-corrupted set presents a great challenge for modern DNNs. Therefore, to investigate learning with noisy labels (LNL) [7, 8, 9, 5] is practically important, which contributes to improvements of model’s generalization on real-world applications.
Figure 1: Existing selection criteria always lead to the imbalanced training set. We train ResNet-18 on noisy CIFAR-100 with two common selection strategies, including small-loss selection [10] and fluctuation-based noise filtering [11]. Class-level selected samples in the last epoch are counted.
Sample selection [10, 12, 11], a prevailing strategy for LNL, achieves considerable performance in mitigating the effects of noisy labels [13] by carefully selecting clean samples from the label-corrupted training set. The performance of sample selection approaches is largely decided by the selection criteria, which can be roughly categorized into two sides: 1) small-loss based strategies [10, 8, 14, 12, 15, 16], which are motivated by the memorization effect [17] that DNNs learn simple patterns shared by majority examples before fitting the noise. Hence, the samples with small losses in the early learning stage can normally be taken as clean samples. 2) fluctuation based strategies [18, 9, 11, 19], which are motivated by the observation that DNNs easily give inconsistent prediction results for noisy samples [11]. These methods normally consider an instance incorrectly labeled if its prediction results exhibit alterations within two consecutive training rounds or during a specified subsequent period.
Despite the validated effectiveness, there is a common view shared by existing sample selection methods that the sample selection processes are only influenced by a training bias, i.e., the accumulated error. In this paper, we for the first time show that by extensive experiments there exists another type of bias, i.e., the data bias, which is mainly caused by imbalanced data distribution. To be concrete, we split the bias in a selection-based learning framework into the training bias, which inherently exists in a self-training manner, and the data bias, which indicates the selected set tends to be class-imbalanced (see Figure 1). Therefore, the model is usually trained on class-imbalanced biased sets, resulting in sub-optimal results. Meanwhile, our experiments also provide some nontrivial findings, including 1) a corrupted selection tendency (see Figure 1(a)). The manner of self-training would exacerbate the imbalanced ratio in the training process. 2) class-discriminative selection performance (see Figure 1(b)). Although the overall selection performance is gradually improved as the training process proceeds, the performance on tail classes exhibits obvious degeneration. The above observations motivate us to consider a novel learning framework that can simultaneously deal with both training bias and data bias.
To this end, we propose a noIse-Tolerant Expert Model (ITEM), consisting of a noise-robust multi-experts structure and a mixed sampling strategy. Firstly, motivated by the idea of ensemble learning [20, 21], we design a robust network architecture that integrates the classifier with multiple experts to mitigate the training bias. Specifically, compared with the prevailing double-branch network [10, 12] where one network feeds selected samples to the other network, we independently conduct classifier learning and sample selection on different output layers. Therefore, incorrect selection results yielded by expert layers would not directly degenerate the classifier in subsequent learning, mitigating the accumulated error to a large degree. Meanwhile, compared with the double-branch network, our network is resource-friendly thanks to fewer trainable parameters and achieves better performance by combining multiple experts. Secondly, to confront the data bias, we propose a mixed sampling training strategy based on weighted data sampling. By calculating the quantity per class in the selected set, we obtain weights of different classes and further reverse weights via a mapping function for assigning larger weights to tail classes when sampling. Eventually, by using the mixup operation [22], our network is trained on a mixed mini-batch that combines two mini-batches respectively yielded by these two weighted samplers. This strategy fulfills class-balanced learning while avoiding sparse representations caused by the sampling strategy.
Our contribution can be summarized as follows:
  • We propose a noise-robust network architecture. Our multi-experts network is resource-friendly and exhibits improved robustness compared with the double-branch network. It shows significant potential in becoming the backbone to tackle noisy labels.
  • We propose a mixed sampling strategy, enabling the model to learn from tail classes via weighted samplers while avoiding sparse representations.
  • ITEM achieves significant performance on both synthetic and real-world noisy datasets, ensuring the effectiveness of our proposal on practical applications.
We also provide abundant analyses, verifying the effectiveness of each component in ITEM.
2 Related Work
Sample selection for LNL. Previous sample selection methods normally exploit the memorization effect of DNNs, i.e., DNNs first memorize training data with clean labels and then those with noisy labels [10, 23]. The small-loss criterion [10] is a typical method stemming from the memorization effect, which splits the noisy training set via a loss threshold. The samples with small losses are regarded as clean samples. Moreover, some researchers modified the small-loss criterion and proposed to select by leveraging the idea of disagreement [14], a regularization term (the Jensen-Shannon Divergence) [8], and a Mixture Gaussian Model for fitting the loss distribution [12], respectively. Besides, to mitigate the accumulated error caused by sample selection, current methods always resort to a double-branch network structure that ensembles selection results from two networks [10, 12, 8, 14, 24]. Besides, some studies [9, 18] introduce the uncertainty of loss to identify the noisy labels. i.e., predictions of noisy instances easily change and thus yield uncertain losses. A recent work [11] points out that hard samples are easily mixed with noisy samples and further discarded by the small-loss criterion. To solve this problem, Wei et al. [11] proposed a fluctuation-based sample selection method that considers the samples with inconsistent prediction results as incorrectly labeled data.
Long-tailed classification. Current long-tailed classification methods can be roughly categorized into two strategies, 1) customize the data sampler to modify sampling frequency for different classes (including over-sampling the tail-class samples and under-sampling the head-class samples) [25, 26, 27, 28], and 2) generate class-wise weights of training losses for improving the importance of tailed classes in terms of empirical risk minimization [29, 25, 28, 30, 31, 32]. However, both strategies result in underfitting in the representation of head classes, i.e., the model eventually learns sparse feature representations for head classes [33, 34]. To overcome this issue, some studies [35, 36] leverage the knowledge underlying head classes to promote the learning of tail classes. Cui et al. [37] designs a re-balance loss by computing the number of instances from each class and assigning different importance weights. Tang et al. [38] leverages causal learning and disentangles feature learning via backdoor adjustment.
(a) A corrupted selection tendency
(b) Class-level selection performance
Figure 2: We train ResNet-18 on Sym. 60% CIFAR-100 with small-loss [10] and fluctuation-based noise filtering [11]. (a) Results on small-loss. Left: the class distribution of three training stages. Right: dynamic numbers of four representative categories. (b) Results on two selection criteria. Visualization of the class-level selection performance $\text{F}_{k}$ in 100- and 200-th epoch. The average F-score among all classes is reported. Although the performance increases, the change is different for varying classes.
Mixture-of-experts. MoE, which belongs to a combination method, has great potential to improve performance in machine learning [39]. Earlier works on MoE mainly focus on dividing the problem space between different experts. Concretely, combining classifiers exhibits promising performance in classification tasks [40, 41], especially for the problem that involves a limited number of patterns, high-dimensional feature sets, and highly overlapped classes [42, 43]. Nowadays, MoE has been applied to several fields, such as long-tailed classification [44, 45] and nature languages processing (NLP) [46, 47].
In this paper, we do not focus on the selection criterion but rather propose a new robust network architecture. Generally, our network can be widely applied to current sample selection methods and further improve their performance.
3 Background
In this section, we first introduce the formal problem setting of learning with noisy labels (LNL) and its learning goal. Then, we conduct experiments to deeply analyze where the bias of sample selection in LNL comes from.
Assume $\mathcal{X}$ is the feature space and $\mathcal{Y}=\{1,2,\ldots,K\}$ is the label space. Suppose the training set is denoted by $D_{N}=\{({\bm{x}}_{i},y_{i})\}_{i\in[N]}$, where $[N]=\{1,2,\ldots,N\}$ is the set of indices. Since the annotator may give wrong labels in practice [48, 49], the learner thus can only observe a label-corrupted set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$.
As a prevailing strategy for LNL, sample selection [11, 50, 12, 9] aims to progressively select a reliable subset ${\tilde{D}_{M}}\subseteq{\tilde{D}_{N}}$ ($M<N$) and feed ${\tilde{D}_{M}}$ to the classifier for training. We introduce $v_{i}\in\{0,1\}$ to indicate whether the $i$-th instance is selected ($v_{i}=1$) or not ($v_{i}=0$). The performance of selection can be reflected by the deviation between the actual selected samples and the total clean samples, which can be measured by the F-score F, where $\text{F}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$, P and R denote selection precision and recall, respectively. In this paper, we individually calculate each class’s F-score for further analyses, i.e., $\text{F}_{k}$=$\frac{2\cdot\text{P}_{k}\cdot\text{R}_{k}}{\text{P}_{k}+\text{R}_{k}}$, $\text{P}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,\tilde{y}_{i}=k)}$, and $\text{R}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(y_{i}=\tilde{y}_{i}=k)}$.
By analyzing the selection performance under different training conditions, we have several nontrivial findings.
  • Current selection criteria easily result in an imbalanced selection subset. As shown in Figure 1, we can observe that both the two adopted selection strategies incur the imbalanced data distribution under two noisy conditions. The reason is that both the two selection criteria rely on model performance. However, the model has different capacities for different classes. For those classes with indistinguishable characteristics (i.e., tail classes in the figure), the model tends to produce a large loss or inconsistent results and further discards those samples.
  • Sample selection, conducted as self-training, will exacerbate the imbalanced ratio during the training process. As shown in Figure 1(a), the imbalanced ratio of the selected set increases as the training process proceeds, i.e., the number of selected instances increases in head classes and decreases in tail classes (see the right part of Figure 1(a)). The intuitive reason is that the model’s performance on tail classes hardly improves due to the limited number of available samples for training, which further degrades the effectiveness of selection criteria on these classes.
  • The selection performance is inherently influenced by the imbalanced class distribution. As shown in Figure 1(b), the selection performance of both two selection criteria increases in head classes but decreases in tail classes. The reason is that the selected errors are relatively small for categories with higher F-scores, hence the training of these classes would further improve the model’s accuracy and selection performance. However, for tail classes with lower F-score, the self-training mechanism would further increase the selection error, and thus fail to correctly select instances in the follow-up phase.
Figure 3: Comparisons of various networks. (a) Typical classification network, which consists of a feature extractor and a classifier layer. (b) Mixture-of-experts (MoE) [51], a set of experts jointly gives the predicted label for the input. (c) Double-branch robust structure, the network is trained on a selected set which is considered as the clean by another network. (d) Ours, our framework consists of a classifier layer and a set of expert layers.
In LNL, we are given a label-corrupted training set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ of $N$ samples. A classifier with learnable parameters $f_{\bm{\theta}}$ is a function that maps from the input space $\mathcal{X}$ to the label space $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$. In multi-class classification, we always update the parameter $\bm{\theta}$ by minimizing the following empirical risk:
$\hat{R}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i}),$ (1)
where $\mathcal{L}(\cdot)$ is the given loss function, e.g., Softmax Cross-Entropy (CE) loss. In sample selection strategy, previous works mainly focus on designing high-effectively selection criteria. A universal training objective on the selected clean set can be written as
$\small\hat{R}_{\text{clean}}(f)$ $=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(f,\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]},\quad\text{where}\,\,\textsf{Criterion}(\cdot)=\left\{\begin{array}[]{lr}1,\,\,\text{If selected},&\\ 0.&\end{array}\right.$ (4)
We empirically show that directly optimizing ${\hat{R}}_{\text{clean}}(f)$ with an existing selection criterion is always biased (in Sec. 3), further resulting in inferior performance. For this, we propose to fulfill debiased learning by handling both the training and data bias. First, to solve the training bias, we design a novel network architecture, called noIse-Tolreant Expert Model (ITEM). Compared with the widely applied double-branch network, our proposal exhibits greater robustness to noisy labels. Second, to solve the data bias, we propose a mixed sampling strategy that can mitigate the side-effect caused by the class-imbalanced set $\tilde{D}_{\text{clean}}$ while avoiding sparse representations.
3.1 ITEM: noIse-Tolerant Expert Model
The training bias of sample selection stems from the way of a self-training manner. Previous works [10, 12] always resort to the double-branch network to confront this problem. Concretely, the network is trained on the clean set that is selected by the other network. Motivated by Mixture-of-experts [41, 51], we propose a robust architecture that integrates multiple experts into the classifier, which independently conducts classifier learning and sample selection in a single network.
Specifically, compared with a normal classification network with a classifier layer $f$, our network ITEM contains $f$ and a set of expert layers $\{g^{1},...,g^{m}\}$ with size $m$. For robust selection, we propose to conduct existing criteria on expert layers. Therefore, the training objective in our framework can be summarized as
$\small\hat{R}_{\text{clean}}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\{g^{1},...,g^{m}\},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$
Intuitively, the independent selection phase avoids the classifier layer selecting noisy samples and then updates its parameters on these data, which contributes to mitigating the training bias. We show the comparison of varying networks in Figure 3. Compared with the prevailing double-branch network, which is proved to be noise-robust in previous works [12, 10], our architecture contains two merits, including 1) resource-friendly training, in contrast to the feature extractor’s huge parameters, the last layer’s parameters for classification are significantly fewer. For example, the parameters of the classifier (or the expert) layer are less than 0.01 million, while the total parameters of a single ResNet-18 for CIFAR-10 are roughly 11.1 million. Thus, our network is more friendly to GPU and computer memory. 2) Greater performance, thanks to the property of fewer parameters, our model can integrate more experts, benefiting both selection and test accuracy by ensembling the results.
Figure 4: The mapping function $\mathcal{S}^{\beta}(\cdot)$ with varying values of $\beta$. Note that $\mathcal{S}(\cdot)$ degrades into a transverse line when $\beta=1$.
Remark. How to update parameters of the expert layer? All selection criteria depend on the model’s performance, including both small-loss family and fluctuation-based methods. Therefore, leveraging expert layers for selecting without updating parameters easily results in poor selection, even leading to a collapse. To fulfill accurate selection on expert layers, we propose a stochastic updating strategy, which randomly assigns a layer from the set of $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$ as the classifier layer $f^{c}$ and the rest are expert layers. Hence, Eq. (3.1) can be rewritten as
$\small\hat{R}_{\text{clean}}(f^{c})=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\mathbbm{F}\backslash f^{c},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f^{c}(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (5)
Since sufficient training processes, we consider that each layer in ITEM contains remarkable performance on selection and prediction.
3.2 Mixed Sampling
To overcome the data bias, we propose a Mixed Sampling strategy to simultaneously learn from head and tail classes in each training iteration. Specifically, supposing a clean set ${\tilde{D}}_{\text{clean}}$ is selected from the ${\tilde{D}_{N}}$, where ${\tilde{D}}_{\text{clean}}\subseteq{\tilde{D}_{N}}$. By calculating the number of instances from each class in ${\tilde{D}}_{\text{clean}}$ and conducting L1 normalization, a vector of class weights $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ can be obtained. The weight of class $k$ is written as
$\bm{w}_{k}=\texttt{Normalize}(S_{k})=u_{k}\Big{/}\sum\nolimits_{i=1}^{K}u_{i},$ (6)
where $u_{i}$ denotes the quantity of $i$-th class in ${\tilde{D}}_{\text{clean}}$.
Based on the weight vector $\bm{v}$, we further propose a reversed weight vector that assigns a larger weight to tailed classes when sampling. Concretely, we construct a mapping function based on the probability density function (PDF) of a Beta distribution. Normally, the Beta distribution’s PDF is controlled by two parameters $\alpha,\beta$. To satisfy both 1) the input interval of $\bm{w}_{k}$ belongs to $\{0,1\}$ and 2) the function contains the monotonically decreasing property, we thus fix $\alpha=1$ and adjust $\beta$ to achieve different slopes. Given a weight $\bm{w}$, the reversed weight ${\bm{\tilde{w}}}$ can be obtained by our mapping function, which is written as
${\bm{\tilde{w}}}=\mathcal{S}^{\beta}(\bm{w})=\frac{1}{\text{B}(1,\beta)}\,(1-\bm{w})^{\beta-1},$ (7)
[…]
Algorithm 1 ITEM
Input: Noisy training set $\tilde{D}_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyper-parameter $\beta$.
1:  Initialize our network with one classifier layer and $m$ expert layers $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$.
2:  for $t$ = 1, …, $T_{w}$ do
3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
4:  end for
5:  for ${t}$ = 1, …, $T$ do
6:     Select the clean set ${\tilde{D}}_{\mathrm{clean}}$ from $D_{N}$ based on $\textsf{Criterion}(\mathbbm{F})$.
7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\mathrm{clean}}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
9:     while $i$ = 1, …, iterations do
10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{v}$.
11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{\tilde{v}}$.
12:        Randomly sample $f^{c}$ from $\mathbbm{F}$.
13:        Update the network parameters by minimizing $\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)}$.
14:     end while
15:  end for
Training on the tail-focused training batch directly would result in the sparse representation of head classes [52]. To avoid this problem, we separately conduct two times weighted sampling (according to the weighted and reversed weighted vector ${\bm{v,\tilde{v}}}$) in data loader phase. Then, we have two training mini-batches $B_{\bm{v}}$ and $B_{{\bm{\tilde{v}}}}$, which focus on head and tail classes, respectively. By leveraging MixUp strategy [22], the model can learn a more representative feature extractor on the mixed data $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$.
Table 1: Test accuracy (mean±$\pm$±std) of methods using ResNet-18/34 on CIFAR-10/100. Note that $\dagger$,$\ddagger$ , ‡ and $\sharp$ denote three selection criteria, i.e., small-loss detection with loss threshold [10], small-loss detection with Gaussian Mixture Model [12], and fluctuation-based noise filtering [11]. Bold values denote the best the second best performance.
Methods Sym. 20% Sym. 40% Inst. 20% Inst. 40% Avg.
CIFAR-10 Cross-Entropy [53] 85.00 $\pm$ 0.43% 79.59 $\pm$ 1.31% 85.92 $\pm$ 1.09% 79.91 $\pm$ 1.41% 82.61
MentorNet [54] 80.49 $\pm$ 0.11% 77.48 $\pm$ 3.45% 79.12 $\pm$ 0.42% 70.27 $\pm$ 1.52% 76.84
JoCoR [8] 88.69 $\pm$ 0.19% 85.44 $\pm$ 0.29% 87.31 $\pm$ 0.27% 82.49 $\pm$ 0.57% 85.98
Joint Optim [55] 89.70 $\pm$ 0.36% 87.79 $\pm$ 0.20% 89.69 $\pm$ 0.42% 82.62 $\pm$ 0.57% 87.45
CDR [23] 89.68 $\pm$ 0.38% 86.13 $\pm$ 0.44% 90.24 $\pm$ 0.39% 83.07 $\pm$ 1.33% 87.28
Me-Momentum [16] 91.44 $\pm$ 0.33% 88.39 $\pm$ 0.34% 90.86 $\pm$ 0.21% 86.66 $\pm$ 0.91% 89.34
PES [56] 92.38 $\pm$ 0.41% 87.45 $\pm$ 0.34% 92.69 $\pm$ 0.42% 89.73 $\pm$ 0.51% 90.56
Late Stopping [19] 91.06 $\pm$ 0.22% 88.92 $\pm$ 0.38% 91.08 $\pm$ 0.23% 87.41 $\pm$ 0.38% 89.62
$\dagger$ Co-teaching [10] 87.16 $\pm$ 0.52% 83.59 $\pm$ 0.28% 86.54 $\pm$ 0.11% 80.98 $\pm$ 0.39% 84.56
$\dagger$ ITEM (Ours) 93.79 $\pm$ 0.14% 90.83 $\pm$ 0.19% 93.52 $\pm$ 0.14% 91.09 $\pm$ 0.18% 92.31
$\ddagger$ ITEM (Ours) 95.01 $\pm$ 0.21% 93.10 $\pm$ 0.20% 95.18 $\pm$ 0.19% 93.65 $\pm$ 0.12% 94.24
$\sharp$ SFT [11] 92.57 $\pm$ 0.32% 89.54 $\pm$ 0.27% 91.41 $\pm$ 0.32% 89.97 $\pm$ 0.49% 90.87
$\sharp$ ITEM (Ours) 95.26 $\pm$ 0.23% 92.81 $\pm$ 0.20% 95.80 $\pm$ 0.18% 93.13 $\pm$ 0.29% 94.25
CIFAR-100 Cross-Entropy [53] 57.59 $\pm$ 2.55% 45.74 $\pm$ 2.61% 59.85 $\pm$ 1.56% 43.74 $\pm$ 1.54% 51.73
MentorNet [54] 52.11 $\pm$ 0.10% 35.12 $\pm$ 1.13% 51.73 $\pm$ 0.17% 40.90 $\pm$ 0.45% 44.96
JoCoR [8] 64.17 $\pm$ 0.19% 55.97 $\pm$ 0.46% 61.98 $\pm$ 0.39% 50.59 $\pm$ 0.71% 58.17
Joint Optim [55] 64.55 $\pm$ 0.38% 57.97 $\pm$ 0.67% 65.15 $\pm$ 0.31% 55.57 $\pm$ 0.41% 60.81
CDR [23] 66.52 $\pm$ 0.24% 60.18 $\pm$ 0.22% 67.06 $\pm$ 0.50% 56.86 $\pm$ 0.62% 62.65
Me-Momentum [16] 68.03 $\pm$ 0.53% 63.48 $\pm$ 0.72% 68.11 $\pm$ 0.57% 58.38 $\pm$ 1.28% 64.50
PES [56] 68.89 $\pm$ 0.41% 64.90 $\pm$ 0.57% 70.49 $\pm$ 0.72% 65.68 $\pm$ 0.44% 67.49
Late Stopping [19] 68.67 $\pm$ 0.67% 64.10 $\pm$ 0.40% 68.59 $\pm$ 0.70% 59.28 $\pm$ 0.46% 65.16
$\dagger$ Co-teaching [10] 59.28 $\pm$ 0.47% 51.60 $\pm$ 0.49% 57.24 $\pm$ 0.69% 45.69 $\pm$ 0.99% 53.45
$\dagger$ ITEM (Ours) 72.42 $\pm$ 0.17% 71.96 $\pm$ 0.24% 73.61 $\pm$ 0.16% 69.90 $\pm$ 0.30% 71.97
$\ddagger$ ITEM (Ours) 78.20 $\pm$ 0.09% 75.27 $\pm$ 0.20% 77.91 $\pm$ 0.14% 70.69 $\pm$ 0.31% 75.51
$\sharp$ SFT [11] 71.98 $\pm$ 0.26% 69.72 $\pm$ 0.31% 71.83 $\pm$ 0.42% 69.91 $\pm$ 0.54% 70.86
$\sharp$ ITEM (Ours) 77.19 $\pm$ 0.13% 74.90 $\pm$ 0.24% 76.91 $\pm$ 0.23% 71.44 $\pm$ 0.29% 75.11
3.3 Practical Training
Despite the training objective of selection-based methods being summarized as a joint framework (in Eq. (5)), the training procedure is practically decoupled into two stages that are iteratively proceeding, 1) select clean samples via the preset criterion and 2) train the network on the selected set. Our training manner also obeys this framework.
Specifically, supposing a selection criterion $\textsf{Criterion}(\cdot)$ such as small-loss [10] or fluctuation-based noise filtering [11] and a network with $m$ size expert layer. In each training epoch, the clean set ${\tilde{D}}_{\text{clean}}$ is selected. To calculate the number of each class in ${\tilde{D}}_{\text{clean}}$, we have a weighted vector $\bm{v}$. Then, the reversed vector $\bm{\tilde{v}}$ is obtained via the mapping function (in Eq. (7)). In each training iteration, we respectively sample a head-focused batch $B{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ and a tail-focused batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\text{clean}}$ according to ${\bm{\tilde{v}}}$, where $b$ denote the size of the mini-batch. By randomly selecting a layer from $\mathbbm{F}$ as the classifier layer $f^{c}$, the parameter’s update can be achieved by minimizing the following training loss
$L^{\mathrm{train}}=\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)},$ (8)
[…]
4 Experiments
Datasets. We assess the performance of ITEM on two noise-synthetic datasets CIFAR-10 and CIFAR-100 [48], and two real-world noisy datasets CIFAR-10N [57] and Clothing-1M [58]. To manually generate the noise, we follow the work [8] and construct symmetric and instance-dependent label noise on CIFAR-10 and CIFAR-10. Specifically, given a noise ratio $\rho$, symmetric label noise randomly translates the true label to a wrong label with probability $\rho$, while instance-dependent label noise assumes a corresponding relationship exists in input features [59]. In experiments, we set the $\rho\in\{20\%,40\%\}$. For testing real-world performance, we employ CIFAR-10N, a human-annotated dataset with varying noise conditions, including worst and random 1&2&3. Besides, Clothing-1M [58], a large-scale dataset that is collected from the website and contains one million training data with 14 categories, is also utilized.
Compared methods. On CIFAR-10 & 100, we compare ITEM with prevailing methods that can be roughly divided into three parts, including 1) sample selection methods (small-loss), MentorNet [54], JoCoR [8], and Me-Momentum [16], 2) sample selection methods (fluctuation), SFT [11] and Late Stopping [19], 3) others, Cross-Entropy [53], Joint Optim [55], and PES [56]. Note that all reported results in this paper are collected from SFT and Late Stopping. On the real-world dataset CIFAR-10N, we compare ITEM with Co-teaching+ [14], Peer Loss [60], CAL [61], DivideMix [12], ELR+ [15], and CORES* [59]. The results are collected from the literature [57].
Experimental setup. We keep the convention from [11, 19] and adopt ResNet-18 and ResNet-34 for CIFAR-10 and CIFAR-100, respectively. For all noisy conditions on CIFAR-10 & 100, we leverage an SGD optimizer with the momentum 0.9 and the weight-decay 1e-3 to train our network. The total training epoch is set as 200. The initial learning rate is 0.02 and decayed with the factor 10 at the 100-th and 150-th epoch. For real-world noise, ResNet-34 and a pre-trained ResNet-50 are used for Clothing-1M and CIFAR-10N, respectively. More details can be found in the Table in Appendix C.
Hyperparameter setup. Our framework contains two main hyperparameters, i.e., the number of expert $m$ in the network architecture, and the slope parameter $\beta$ in the mapping function. We keep $\beta=2$ for all experiments and set $m=2,4$ for CIFAR and Clothing-1M, respectively.
4.1 Comparison Results
Results on CIFAR-10 and CIFAR-100. Since our framework is not restricted by different selection criteria, we test our proposal on three types of selection criterion, including small-loss with loss threshold [10], small-loss with Gaussian Mixture Model [12] and fluctuation-based noise filtering [11]. Each trial is repeated five times with different seeds. Mean values with standard deviation are recorded.
[…]
Methods CIFAR-10N
Worst R. 1 R. 2 R. 3
CE 77.69 85.02 86.46 85.16
Co-teaching+ 83.26 89.70 89.47 89.54
Peer Loss 82.00 89.06 88.76 88.57
CAL 85.36 90.93 90.75 90.74
Late Stoping 85.27 - - -
$\ddagger$ ITEM 91.15 95.12 95.03 95.17
DivideMix 92.56 95.16 95.23 95.21
ELR+ 91.09 94.43 94.20 94.34
CORES* 91.66 94.45 94.88 94.74
$\ddagger$ ITEM* 93.14 96.08 96.54 96.71
[…]
Methods Test acc. (%) Methods Test acc. (%)
CE 64.54 Co-teaching 69.21
Forward 69.84 SFT+ 75.08
JoCoR 70.30 DMI 72.46
Joint Optim 72.23 PES (semi) 74.85
DivideMix 74.76 ELR+ 74.81
$\ddagger$ ITEM 74.69 $\ddagger$ ITEM* 75.02
Results on CIFAR-10N. We evaluate the effectiveness of ITEM on a human-annotated noisy dataset CIFAR-10N which contains four noisy conditions, i.e., Worst and Random 1&2&3 [57]. Besides, following previous selection-based approaches [12], we discard labels of unreliable samples and train the model with them by semi-supervised learning (SSL) for greater generalization, which is represented as ITEM*. Appendix A shows the detailed algorithm flowchart for ITEM*.
The results can be found in Table 2. First, our method achieves remarkable performance on both two settings (w/ or w/o SSL). When training without SSL, ITEM outperforms previous methods by a large margin. Compared with Late Stopping, a newly proposed method, ITEM achieves 5.88% improvements on worst-labels. When training with SSL, ITEM is also superior to the previous SOTA method DivideMix. It is worth noting that even training without SSL, ITEM surprisingly gains greater performance than those methods that leverage SSL approaches.
Results on Clothing-1M. Going beyond the small-scale benchmark such as CIFAR-10 & 100, we also test the performance of ITEM and ITEM* on a large-scale dataset Clothing-1M, which contains roughly 38.49% noisy labels. The results are shown in Table 3. We can observe that both ITEM and ITEM* outperform existing methods by a meaningful margin. For example, ITEM* achieves 0.17% improvements compared with the second-best method PES. The result verifies that our method is effective for enhancing model generalization in large-scale real-world scenarios.
[…]
$\xi$ 10
$\rho$ 10% 20% 30% 40% 50%
CIFAR-10 Co-teaching 80.30 78.54 68.71 57.10 46.77
Sel-CL+ 86.47 85.11 84.41 80.35 77.27
RCAL 88.09 86.46 84.58 83.43 80.80
Ours 87.32 86.31 84.91 84.72 81.04
CIFAR-100 Co-teaching 45.61 41.33 36.14 32.08 25.33
Sel-CL+ 55.68 53.52 50.92 47.57 44.86
RCAL 57.50 54.85 51.66 48.91 44.36
Ours 56.10 55.93 53.07 48.29 45.17
Table 5: Ablation study of each component in $\ddagger$ ITEM with varying noise conditions.
Methods CIFAR-10 CIFAR-100 CIFAR-10N
Sym 40% Inst. 40% Sym 40% Inst. 40% Worst
w/o robust multi-expert network (training on normal ResNet) 90.75 91.44 73.01 70.04 89.72
w/o mixed sampling (training on a random sampled batch) 88.77 89.91 71.09 68.54 87.66
w/o mixup (directly training on $B{{\bm{v}}}$ and $B{{\bm{\tilde{v}}}}$) 90.19 90.35 72.50 69.11 89.10
Ours 92.81 93.13 74.90 71.44 91.15
(a) The number of expert layers $m$.
(b) The slope parameter $\beta$.
Figure 5: Sensitivity analyses for hyperparameters selection.
4.2 More Analyses
[…]
Ablation study. Our framework mainly contains two modules, the robust network architecture and the mixed data sampling strategy. We conduct ablation studies on CIFAR benchmarks to evaluate the effectiveness of each component (shown in Table 5). First, compared with typical ResNet architecture, our expert-based network indeed makes an obvious contribution to mitigating the noise. The average improvement among five noisy settings roughly reaches 2%. Second, the mixed sampling strategy also promotes the performance of our network. Note that compared with training on randomly sampled batches, training on two weighted sampled batches gains greater generalization. Thus, each module in our learning framework is effective.
Hyperparameter selection. There are two main hyperparameters $m$ and $\beta$ in our framework. We conduct ablation studies to select optimal values for experiments, where $m\in\{2,3,4,5,6,7,8\}$ and $\beta\in\{1.3,1.8,2,3,5,10\}$. The result is shown in Figure 5. For the number of expert layers, we can see that the best test accuracy increases with a larger value of $m$. When $m=4$, the performance reaches the peak and slightly fluctuates as $m$ increases. Meanwhile, the gap between the peak value and the least value is around 0.6% on CIFAR-10 and 0.9% on CIFAR-100, respectively. For the slope parameter in the mapping function, we observe that the performance of our framework is not sensitive to varying values of $\beta$. Therefore, we select $m=4$ and $\beta=3$ for all experiments.
Figure 6: Comparisons of class-level prediction results on CIFAR-100 with 40% symmetric noise. Note that “CE (clean)" denotes training the model on the completely clean set (50k samples).
Class-balanced learning. We visualize the test accuracy in each class to demonstrate that ITEM achieves relatively balanced performance on varying categories while fulfilling greater generalization. We plot comparison results in Figure 6. First, compared with CE (clean), which is trained on 50k clean samples, ITEM achieves almost unbiased prediction results. Second, compared with existing methods, ITEM takes tail classes better into account since two weighted samplers. Therefore, ITEM obtains greater performance in these categories. The result demonstrates our proposal has considerable potential in fulfilling class-balanced learning while tackling label noise.
5 Conclusion
In this paper, we disclosed the data bias, an implicit bias underlying the sample selection strategy. To solve the training and data bias simultaneously, we proposed ITEM, which integrates a noise-robust multi-experts network and a mixed sampling strategy. First, our structure integrates the classifier with multiple experts and leverages expert layers to conduct selection independently. Compared with the prevailing double-branch network, it exhibits great potential in mitigating the training bias, i.e., the accumulated error. Second, our mixed sampling strategy yields class-aware weights and further conducts weighted sampling. By using the mixup of two weighted mini-batches, we mitigate the issue of sparse representation. The effectiveness of ITEM in real applications is verified on diverse noise types and both synthetic and real-world datasets.
Appendix for Debiased Sample Selection for Combating Noisy Labels
A ITEM with a Semi-Supervised Learning Framework
Considering the large scale of unreliable samples discarded, especially in considerable noise ratios, leveraging these samples through semi-supervised learning is necessary. Therefore, we provide a novel version of ITEM with stronger performance. The algorithm is shown in Algorithm 2. Besides, a strong augmentation CTAugment is used, which refers to [64]. Compared with its original version, the new training framework only adds two operations (shown in Lines 12 and 13). Thus, the code implementation is relatively convenient.
Algorithm 2 ITEM with a semi-supervised framework
Input: The training set $D_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyperparameter $\beta$.
1:  Initialize our network with one classifier layer and $m$ expert layers $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$.
2:  for $t$ = 1, …, $T_{w}$ do
3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
4:  end for
5:  for ${t}$ = 1, …, $T$ do
6:     Split $D_{N}$ into the clean set ${\tilde{D}}_{\rm clean}$ and the unreliable set ${\tilde{D}}_{\rm noise}$ based on $\textsf{Criterion}(\mathbbm{F})$, and discard labels in ${\tilde{D}}_{\rm noise}$.
7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\rm clean}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
9:     while $i$ = 1, …, iterations do
10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{v}$.
11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x}^{\prime}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{\tilde{v}}$.
12:        Draw a mini-batch $B=\{(\bm{x}^{\prime\prime}_{i})\}_{i=1}^{2b}$ from ${\tilde{D}}_{\rm noise}$ without labels.
13:        Generate pseudo-labels, then have $B=\{(\bm{x}^{\prime\prime}_{i},{\tilde{y}^{\prime\prime}}_{i})\}_{i=1}^{2b}$, where ${\tilde{y}^{\prime\prime}}_{i}=\mathbbm{E}_{f\sim\mathbbm{F}}[f(\bm{x}^{\prime\prime}_{i})]$.
14:        Randomly sample $f^{c}$ from $\mathbbm{F}$.
15:        Update network parameters via minimizing $\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(B_{{\bm{v}}}\text{ and }B_{{\bm{\tilde{v}}}},B)\big{)}$.
16:     end while
17:  end for
B More experimental results
B.1 Results on CIFAR-100N
We additionally verify the effectiveness of $\ddagger$ ITEM on another human-annotated dataset, CIFAR-100N, whose noise ratio is roughly 40.2%. We adopt the same settings on CIFAR-100 (see Table 8), and the results are shown in Table 6. Our proposal ITEM achieves state-of-the-art performance on both supervised and semi-supervised frameworks.
w/o semi w/ semi
CE 55.50 $\pm$ 0.66% ELR + 66.72 $\pm$ 0.07%
Co-teaching 60.37 $\pm$ 0.27% CORES* 55.72 $\pm$ 0.42%
SOP+ 67.81 $\pm$ 0.23% DivideMix 71.13 $\pm$ 0.48%
ITEM 69.47 $\pm$ 0.18% ITEM* 72.40 $\pm$ 0.19%
Table 6: Test accuracy (%) on the test set of CIFAR-100N.
B.2 Visualization
We verify the effectiveness of $\ddagger$ ITEM on varying noisy conditions and show the following noise-robust results.
Figure 7: Visualization of class-level selection performance of ITEM on CIFAR-100 with 60% symmetric label noise. As the training process proceeds, ITEM gradually improves the F-score in all classes instead of only in head classes.
(a) Training on randomly sampled mini-batch $B$, the best test accuracy is 88.42%.
(b) Training on $B_{\tilde{v}}$ only, the best test accuracy is 90.17%.
(c) Training on $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$, the best test accuracy is 93.04%.
Figure 8: Visualizing feature representations on the CIFAR-10 test set by T-SNE. The noise setting is 40% symmetric noisy labels.
(a) CIFAR-10 $w/$ Inst. 40%
(b) CIFAR-100 $w/$ Inst. 40%
Figure 9: Visualization of training curves under three settings. Suppose a small-loss criterion with the loss threshold, we compare our methods (Ours) with the proposed multi-expert network and single network on two noise conditions.
  • robust selection processes. As shown in Figure 7, the total selection performance increases as the training processes, while our framework has not degraded the selection performance on tail classes.
  • robust feature representations. We visualize the learned representation of the test set on CIFAR-10. We compare three training manners and show the results in Figure 8. Compared with training on randomly sampled mini-batch and $B_{\tilde{v}}$, our mixed sampling strategy yields a feature extractor with a strong representation. As shown in Figure 7(c), the distance among varying class clusters is closer, reflecting that the classifier learns more robust decision boundaries.
  • robust training curves. We conduct ablation studies to verify the effectiveness of our proposed robust network architecture. The results on two noisy settings are shown in Figure 9. We can see that compared with a single network (a normal ResNet structure), our expert network achieves more robustness. Meanwhile, the mixed sampling strategy also contributes to improving generalization performance.
Methods Stanford Dogs CUB-200-2011 Avg
Sym 40% Asy 30% Sym 40% Asy 30%
CE 51.42 58.08 64.01 56.02 57.38
Label Smooth 70.22 64.99 54.39 56.80 61.60
Conf. Penalty 68.69 64.50 52.40 54.33 59.98
Co-teaching 49.15 50.50 46.57 50.60 49.21
JoCoR 49.62 53.59 52.64 51.70 51.89
CE + SNSCL 75.27 64.49 68.83 61.48 67.52
$\ddagger$ ITEM 74.92 68.19 70.04 66.58 69.93
Table 7: Test accuracy (%) on two fine-grained benchmarks with two noisy settings. The best and the second best performances are highlighted with Bold and underline, respectively.
Table 8: Detailed settings about training procedure of ITEM.
CIFAR-10 CIFAR-100 CIFAR-10N (all conditions) Clothing-1M
Statistic data
 Class number 10 10 10 14
 Training size 50,000 50,000 50,000 1,000,000
 Testing size 10,000 10,000 10,000 10,000
Training procedure
 Network ResNet-18 ResNet-34 Pre-trained ResNet-50
 Batch size 64 100
 Epoch 200 10
 Warmup epoch 10 30 10 1
 Optimizer SGD
 Momentum 0.9
 Learning rate (LR) 0.02 0.001
 Weight decay 1e-3 5e-4
 LR scheduler divided by 10 at [100,150]-th epoch divided by 10 at 5-th epoch
Hyperparameters
 the num. of experts $m$ $m=4$ $m=2$
 the slope parameter $\beta$ $\beta=3$
B.3 Results on Fine-Grained Noisy Settings
The literature [5] pointed out that noisy label’ generation is strongly associated with fine-grained datasets. Following settings in SNSCL [5], we conduct experiments on four noisy fine-grained datasets, including Stanford Dogs [65] and CUB-200-2011 [66]. A ResNet-18 pre-trained on ImageNet with three experts ($m=3$) is utilized. We synthetically construct two noise types, symmetric and asymmetric noise labels.
The results are shown in Table 7. Our method achieves significant generalization performance even when tackling fine-grained noisy classification and consistently outperforms other methods by a large margin. We consider the MixUp operation to be the primary contribution to the advanced performance on fine-grained datasets since it encourages the model to learn from structured data instead of the unstructured noise [67].
C Experimental Settings
We reported the statistical data of four datasets and the whole experimental setup in Table 8. Besides, the data augmentation strategy for different datasets can be divided into two parts.
  • For CIFAR-10, CIFAR-100, and CIFAR-10N, randomly cropping and horizontally flipping are adopted.
  • For Clothing-1M, we first resize the image to 256x256 and then randomly crop it to 224x224. Besides, the horizontally flipping is adopted.
Meanwhile, we detailedly exhibit the noise generation process by the noise transition matrix. As shown in Figure 10, given a noise ratio $\rho$, the noise transition matrix $T$ is represented as $T_{ij}(\bm{x})=P({\tilde{y}}=j|y=i)$, where $T_{ij}$ denotes that the label of the instance transits from a clean label $i$ to a noisy label $j$. The generation process of instance-dependent label noise can be found in the work [23, 11].
Figure 10: Visualization of the transition matrix for symmetric (left) and asymmetric (right) noise on a ten classification task when the noise ratio is $40\%$.
C.1 Settings for Learning with Noisy Labels on Imbalanced Datasets
We obey the same manner of imbalanced noise construction from RCAL [62] and conduct experiments on synthetic CIFAR-10 and CIFAR-100. Concretely, we set the batch size as 64 and run 200 epochs for both CIFAR-10 and CIFAR-100. The learning rate is set as 0.02 and reduced by a factor of 10 at {100, 150}-th epoch. A ResNet-32 network with 4 experts is utilized. Besides, for a fair comparison, we adopt the same augmentation policy SimAug [68] as RCAL.

### v3

@@
-Debiased Sample Selection for Combating Noisy Labels
+Understanding and Mitigating the Bias in Sample Selection for Learning with Noisy Labels
 Abstract
@@
 The sample selection strategy achieves promising performance by selecting a label-reliable subset for model training.
-In this paper, we empirically reveal that existing sample selection methods suffer from both data and training bias that are represented as imbalanced selected sets and accumulation errors in practice, respectively.
+In this paper, we empirically reveal that existing sample selection methods suffer from both data and training bias, which are represented in practice as imbalanced selected sets and accumulation errors.
 However, only the training bias was handled in previous studies.
 To address this limitation, we propose a noIse-Tolerant Expert Model (ITEM) for debiased learning in sample selection.
-Specifically, to mitigate the training bias, we design a robust network architecture that integrates with multiple experts.
-Compared with the prevailing double-branch network, our network exhibits better performance of selection and prediction by ensembling these experts while training with fewer parameters.
-Meanwhile, to mitigate the data bias, we propose a mixed sampling strategy based on two weight-based data samplers.
-By training on the mixture of two class-discriminative mini-batches, the model mitigates the effect of the imbalanced training set while avoiding sparse representations that are easily caused by sampling strategies.
-Extensive experiments and analyses demonstrate the effectiveness of ITEM.
-Our code is available at this url ITEM.
+Specifically, to mitigate the training bias, we design a robust Mixture-of-Expert network that conducts selection and learning on different layers.
+Compared with the prevailing double-branch network, our network performs better on both selection and prediction by ensembling multiple experts while training with fewer parameters.
+Meanwhile, to mitigate the data bias, we propose a weighted sampling strategy that assigns larger sampling weights to classes with smaller frequencies.
+Using MixUp, the model is trained on a mixture of two batches: one sampled by a weighted sampler and the other by a regular sampler, which mitigates the effect of the imbalanced training set while avoiding sparse representations that are easily caused by sampling strategies.
+Extensive experiments on seven noisy benchmarks and analyses demonstrate the effectiveness of ITEM.
+The code is released at https://github.com/1998v7/ITEM.
 1 Introduction
-The remarkable generalization capability of deep neural networks (DNNs) is achieved through training on a large-scale dataset [1, 2].
+The remarkable generalization capability of deep neural networks (DNNs) is achieved through training on a large-scale dataset.
 However, existing training sets are usually collected by online queries [6], crowdsourcing [48], and manual annotations, which could inevitably incur wrong (or noisy) labels [40].
-Since DNNs exhibit vulnerability to such low-quality datasets [6], training on the label-corrupted set presents a great challenge for modern DNNs.
-Therefore, to investigate learning with noisy labels (LNL) [7, 8, 9, 5] is practically important, which contributes to improvements of model’s generalization on real-world applications.
-Figure 1: Existing selection criteria always lead to the imbalanced training set.
-We train ResNet-18 on noisy CIFAR-100 with two common selection strategies, including small-loss selection [10] and fluctuation-based noise filtering [11].
-Class-level selected samples in the last epoch are counted.
-Sample selection [10, 12, 11], a prevailing strategy for LNL, achieves considerable performance in mitigating the effects of noisy labels [13] by carefully selecting clean samples from the label-corrupted training set.
-The performance of sample selection approaches is largely decided by the selection criteria, which can be roughly categorized into two sides: 1) small-loss based strategies [10, 8, 14, 12, 15, 16], which are motivated by the memorization effect [17] that DNNs learn simple patterns shared by majority examples before fitting the noise.
-Hence, the samples with small losses in the early learning stage can normally be taken as clean samples.
-2) fluctuation based strategies [18, 9, 11, 19], which are motivated by the observation that DNNs easily give inconsistent prediction results for noisy samples [11].
-These methods normally consider an instance incorrectly labeled if its prediction results exhibit alterations within two consecutive training rounds or during a specified subsequent period.
-Despite the validated effectiveness, there is a common view shared by existing sample selection methods that the sample selection processes are only influenced by a training bias, i.e., the accumulated error.
-In this paper, we for the first time show that by extensive experiments there exists another type of bias, i.e., the data bias, which is mainly caused by imbalanced data distribution.
-To be concrete, we split the bias in a selection-based learning framework into the training bias, which inherently exists in a self-training manner, and the data bias, which indicates the selected set tends to be class-imbalanced (see Figure 1).
-Therefore, the model is usually trained on class-imbalanced biased sets, resulting in sub-optimal results.
-Meanwhile, our experiments also provide some nontrivial findings, including 1) a corrupted selection tendency (see Figure 1(a)).
-The manner of self-training would exacerbate the imbalanced ratio in the training process.
-2) class-discriminative selection performance (see Figure 1(b)).
-Although the overall selection performance is gradually improved as the training process proceeds, the performance on tail classes exhibits obvious degeneration.
-The above observations motivate us to consider a novel learning framework that can simultaneously deal with both training bias and data bias.
-To this end, we propose a noIse-Tolerant Expert Model (ITEM), consisting of a noise-robust multi-experts structure and a mixed sampling strategy.
-Firstly, motivated by the idea of ensemble learning [20, 21], we design a robust network architecture that integrates the classifier with multiple experts to mitigate the training bias.
-Specifically, compared with the prevailing double-branch network [10, 12] where one network feeds selected samples to the other network, we independently conduct classifier learning and sample selection on different output layers.
-Therefore, incorrect selection results yielded by expert layers would not directly degenerate the classifier in subsequent learning, mitigating the accumulated error to a large degree.
-Meanwhile, compared with the double-branch network, our network is resource-friendly thanks to fewer trainable parameters and achieves better performance by combining multiple experts.
-Secondly, to confront the data bias, we propose a mixed sampling training strategy based on weighted data sampling.
-By calculating the quantity per class in the selected set, we obtain weights of different classes and further reverse weights via a mapping function for assigning larger weights to tail classes when sampling.
-Eventually, by using the mixup operation [22], our network is trained on a mixed mini-batch that combines two mini-batches respectively yielded by these two weighted samplers.
-This strategy fulfills class-balanced learning while avoiding sparse representations caused by the sampling strategy.
-Our contribution can be summarized as follows:
-  • We propose a noise-robust network architecture.
-Our multi-experts network is resource-friendly and exhibits improved robustness compared with the double-branch network.
-It shows significant potential in becoming the backbone to tackle noisy labels.
-  • We propose a mixed sampling strategy, enabling the model to learn from tail classes via weighted samplers while avoiding sparse representations.
-  • ITEM achieves significant performance on both synthetic and real-world noisy datasets, ensuring the effectiveness of our proposal on practical applications.
-We also provide abundant analyses, verifying the effectiveness of each component in ITEM.
-2 Related Work
-Sample selection for LNL.
-Previous sample selection methods normally exploit the memorization effect of DNNs, i.e., DNNs first memorize training data with clean labels and then those with noisy labels [10, 23].
-The small-loss criterion [10] is a typical method stemming from the memorization effect, which splits the noisy training set via a loss threshold.
-The samples with small losses are regarded as clean samples.
-Moreover, some researchers modified the small-loss criterion and proposed to select by leveraging the idea of disagreement [14], a regularization term (the Jensen-Shannon Divergence) [8], and a Mixture Gaussian Model for fitting the loss distribution [12], respectively.
-Besides, to mitigate the accumulated error caused by sample selection, current methods always resort to a double-branch network structure that ensembles selection results from two networks [10, 12, 8, 14, 24].
-Besides, some studies [9, 18] introduce the uncertainty of loss to identify the noisy labels.
-i.e., predictions of noisy instances easily change and thus yield uncertain losses.
-A recent work [11] points out that hard samples are easily mixed with noisy samples and further discarded by the small-loss criterion.
-To solve this problem, Wei et al.
-[11] proposed a fluctuation-based sample selection method that considers the samples with inconsistent prediction results as incorrectly labeled data.
-Long-tailed classification.
-Current long-tailed classification methods can be roughly categorized into two strategies, 1) customize the data sampler to modify sampling frequency for different classes (including over-sampling the tail-class samples and under-sampling the head-class samples) [25, 26, 27, 28], and 2) generate class-wise weights of training losses for improving the importance of tailed classes in terms of empirical risk minimization [29, 25, 28, 30, 31, 32].
-However, both strategies result in underfitting in the representation of head classes, i.e., the model eventually learns sparse feature representations for head classes [33, 34].
-To overcome this issue, some studies [35, 36] leverage the knowledge underlying head classes to promote the learning of tail classes.
-Cui et al.
-[37] designs a re-balance loss by computing the number of instances from each class and assigning different importance weights.
-Tang et al.
-[38] leverages causal learning and disentangles feature learning via backdoor adjustment.
-(a) A corrupted selection tendency
-(b) Class-level selection performance
+Since DNNs exhibit vulnerability to such low-quality training sets [52], training on the label-corrupted set presents a great challenge for the modern machine-learning community.
+To mitigate the adverse effect brought by noisy samples, learning with noisy labels (LNL) [37, 44] is important, contributing to improvements of the model’s generalization on practical applications.
+Figure 1: Existing selection criteria always lead to an imbalanced training set, termed as the data bias.
+A ResNet-34 is trained on CIFAR-100N.
+We visualize the class distribution of the selected set, given three typical selection criteria.
+The quantity of class-level samples in the last epoch is counted, while the index of classes is sorted.
+More results can be found in Appendix A.
 Figure 2: We train ResNet-18 on Sym.
@@
 (b) Results on two selection criteria.
-Visualization of the class-level selection performance $\text{F}_{k}$ in 100- and 200-th epoch.
-The average F-score among all classes is reported.
-Although the performance increases, the change is different for varying classes.
-Mixture-of-experts.
-MoE, which belongs to a combination method, has great potential to improve performance in machine learning [39].
-Earlier works on MoE mainly focus on dividing the problem space between different experts.
-Concretely, combining classifiers exhibits promising performance in classification tasks [40, 41], especially for the problem that involves a limited number of patterns, high-dimensional feature sets, and highly overlapped classes [42, 43].
-Nowadays, MoE has been applied to several fields, such as long-tailed classification [44, 45] and nature languages processing (NLP) [46, 47].
-In this paper, we do not focus on the selection criterion but rather propose a new robust network architecture.
-Generally, our network can be widely applied to current sample selection methods and further improve their performance.
-3 Background
+Visualization of the class-level selection performance (F-score) $\text{F}_{k}$ in 100- and 200-th epoch.
+Sample selection [11, 21, 39], a prevailing strategy for LNL, achieves considerable performance in mitigating the effects of noisy labels [47] by carefully selecting clean samples from the label-corrupted training set.
+The performance of sample selection approaches is largely decided by the selection criteria, which can be roughly categorized into two sides: 1) small-loss based strategies [37, 11, 21, 50, 26, 3], which are motivated by the memorization effect [2] that DNNs learn simple patterns shared by majority examples before fitting the noise.
+Hence, the samples with small losses in the early learning stage can normally be taken as clean samples.
+2) fluctuation based strategies [56, 39, 51, 44], which are motivated by the observation that DNNs easily give inconsistent prediction results for noisy samples.
+These methods [39, 51] normally consider an instance incorrectly labeled if its prediction results exhibit alterations within two consecutive training rounds or during a specified subsequent period.
+Despite the validated effectiveness, there is a common view shared by existing sample selection methods that the sample selection processes are only influenced by a training bias, i.e., the accumulated error.
+In this paper, we, for the first time, show that by extensive experiments, there exists another type of bias, i.e., the data bias, which is mainly caused by imbalanced data distribution.
+To be concrete, we split the bias in a selection-based learning framework into the training bias, which inherently exists in a self-training manner, and the data bias, which indicates the selected set tends to be class-imbalanced (see Figure 1).
+Moreover, we customize a class-level F-score and do extensive experiments in the next section, providing three nontrivial findings that reveal the adverse effect of the data bias.
+These observations motivate us to consider a novel selection-based learning framework that can simultaneously address training and data bias.
+To this end, we propose a noIse-Tolerant Expert Model (ITEM), consisting of a noise-robust multi-experts structure and a weighted sampling strategy.
+  • Firstly, motivated by the idea of ensemble learning [9] and the Mixture-of-Expert model [16], we design a robust network structure integrating the classifier with multiple experts.
+Our network is more robust compared with existing robust networks like double-branch network [11] for the following reasons, 1) in our structure, the classifier learning and selection are disentangled, with experts selecting clean samples for the classifier’s training, naturally mitigating the issue of error accumulation.
+2) Ensembling selection results from different experts reduces harmful interference from incompatible patterns (e.g., poor selection made by individual experts), leading to a cleaner selected set.
+  • Secondly, to mitigate the implicit data bias during selection, we propose a weighted resampling strategy named mixed sampling.
+By calculating the quantity of each class in the selected set, we assign larger weights to tail classes through a mapping function during sampling.
+Finanlly, introducing the MixUp [53], our network is trained on a mixed batch that combines two batches based on a regular and a weighted sampler, fulfilling class-balanced learning while avoiding the issue of sparse representations.
+The promising performance of ITEM is verified on seven noisy benchmarks. Extensive ablation studies and analyses also demonstrate the effectiveness of each component.
+2 Understanding the Bias in Sample Selection
 In this section, we first introduce the formal problem setting of learning with noisy labels (LNL) and its learning goal.
@@
 Assume $\mathcal{X}$ is the feature space and $\mathcal{Y}=\{1,2,\ldots,K\}$ is the label space.
-Suppose the training set is denoted by $D_{N}=\{({\bm{x}}_{i},y_{i})\}_{i\in[N]}$, where $[N]=\{1,2,\ldots,N\}$ is the set of indices.
-Since the annotator may give wrong labels in practice [48, 49], the learner thus can only observe a label-corrupted set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$.
-As a prevailing strategy for LNL, sample selection [11, 50, 12, 9] aims to progressively select a reliable subset ${\tilde{D}_{M}}\subseteq{\tilde{D}_{N}}$ ($M<N$) and feed ${\tilde{D}_{M}}$ to the classifier for training.
+Suppose the training set is denoted by ${\tilde{D}}=\{({\bm{x}}_{i},y_{i})\}_{i\in[N]}$, where $[N]=\{1,2,\ldots,N\}$ is the set of indices.
+Since the annotator may give wrong labels in practice [18], the learner thus can only observe a label-corrupted set ${D_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ with noisy labels.
+As a prevailing strategy for LNL, sample selection [39, 24, 21, 44] aims to progressively select a reliable subset ${D_{M}}\subseteq{D_{N}}$ ($M<N$) and feed ${D_{M}}$ to the classifier for training.
 We introduce $v_{i}\in\{0,1\}$ to indicate whether the $i$-th instance is selected ($v_{i}=1$) or not ($v_{i}=0$).
-The performance of selection can be reflected by the deviation between the actual selected samples and the total clean samples, which can be measured by the F-score F, where $\text{F}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$, P and R denote selection precision and recall, respectively.
-In this paper, we individually calculate each class’s F-score for further analyses, i.e., $\text{F}_{k}$=$\frac{2\cdot\text{P}_{k}\cdot\text{R}_{k}}{\text{P}_{k}+\text{R}_{k}}$, $\text{P}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,\tilde{y}_{i}=k)}$, and $\text{R}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(y_{i}=\tilde{y}_{i}=k)}$.
+Generally, the performance of a selection-based method can be reflected by the deviation between the actual selected samples and the total clean samples, which can be measured by the F-score F, where $\text{F}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$, P and R denote selection precision and recall, respectively.
+In this paper, we individually calculate each class’s F-score for further analysis.
+Specifically, the class-level F-score for class $k$ is written as
+Fk=2PkRkPk+Rk,where
+[…]
+[…]
+[…]
+[…]
+Figure 3: Comparisons of different architectures.
+(a) Typical classification network, which consists of a feature extractor and a classifier layer.
+(b) Mixture-of-experts (MoE) [29], a set of experts jointly gives the predicted label for the input.
+(c) Double-branch robust structure, the network is trained on a selected set that is considered clean by another network.
+(d) Ours, a mixture-of-experts module, is integrated into the classification network, which works for robust selection as well as prediction ensemble.
 By analyzing the selection performance under different training conditions, we have several nontrivial findings.
   • Current selection criteria easily result in an imbalanced selection subset.
-As shown in Figure 1, we can observe that both the two adopted selection strategies incur the imbalanced data distribution under two noisy conditions.
-The reason is that both the two selection criteria rely on model performance.
+As shown in Figure 1, we can see that both of the two adopted selection strategies incur the imbalanced data distribution under two noisy conditions.
+The reason is that both of the two selection criteria rely on model performance.
 However, the model has different capacities for different classes.
-For those classes with indistinguishable characteristics (i.e., tail classes in the figure), the model tends to produce a large loss or inconsistent results and further discards those samples.
+For those classes with indistinguishable characteristics (i.e., rare classes), the model tends to produce a large loss or inconsistent results and further discards those samples.
   • Sample selection, conducted as self-training, will exacerbate the imbalanced ratio during the training process.
-As shown in Figure 1(a), the imbalanced ratio of the selected set increases as the training process proceeds, i.e., the number of selected instances increases in head classes and decreases in tail classes (see the right part of Figure 1(a)).
+As shown in Figure 2 (a), the imbalanced ratio of the selected set increases as the training process proceeds, i.e., the number of selected instances increases in head classes and decreases in tail classes (see the right part of Figure 2).
 The intuitive reason is that the model’s performance on tail classes hardly improves due to the limited number of available samples for training, which further degrades the effectiveness of selection criteria on these classes.
   • The selection performance is inherently influenced by the imbalanced class distribution.
-As shown in Figure 1(b), the selection performance of both two selection criteria increases in head classes but decreases in tail classes.
-The reason is that the selected errors are relatively small for categories with higher F-scores, hence the training of these classes would further improve the model’s accuracy and selection performance.
-However, for tail classes with lower F-score, the self-training mechanism would further increase the selection error, and thus fail to correctly select instances in the follow-up phase.
-Figure 3: Comparisons of various networks.
-(a) Typical classification network, which consists of a feature extractor and a classifier layer.
-(b) Mixture-of-experts (MoE) [51], a set of experts jointly gives the predicted label for the input.
-(c) Double-branch robust structure, the network is trained on a selected set which is considered as the clean by another network.
-(d) Ours, our framework consists of a classifier layer and a set of expert layers.
-In LNL, we are given a label-corrupted training set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ of $N$ samples.
+As shown in Figure 2 (b), the selection performance of both selection criteria increases in head classes but decreases in tail classes.
+The reason is that the selected errors are relatively small for categories with higher F-scores.
+Hence, the training of these classes would further improve the model’s accuracy and selection performance.
+However, for tail classes with lower F-score, the self-training mechanism would further increase the selection error and thus fail to select instances in the follow-up phase correctly.
+Based on these observations, we can deduce that the classifier under a sample selection framework implicitly suffers from a data bias, which is, however, rarely studied in previous literature.
+Motivated by this, we aim to design a model that confronts not only accumulated error but also the imbalanced selected set.
+3 Methodology
+Preliminaries.
+In LNL, we are given a label-corrupted training set ${D_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ of $N$ samples.
 A classifier with learnable parameters $f_{\bm{\theta}}$ is a function that maps from the input space $\mathcal{X}$ to the label space $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$.
@@
 In sample selection strategy, previous works mainly focus on designing high-effectively selection criteria.
-A universal training objective on the selected clean set can be written as
-$\small\hat{R}_{\text{clean}}(f)$ $=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(f,\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]},\quad\text{where}\,\,\textsf{Criterion}(\cdot)=\left\{\begin{array}[]{lr}1,\,\,\text{If selected},&\\ 0.&\end{array}\right.$ (4)
-We empirically show that directly optimizing ${\hat{R}}_{\text{clean}}(f)$ with an existing selection criterion is always biased (in Sec.
-3), further resulting in inferior performance.
-For this, we propose to fulfill debiased learning by handling both the training and data bias.
-First, to solve the training bias, we design a novel network architecture, called noIse-Tolreant Expert Model (ITEM).
+Here, we give a universal training objective on the selected clean set, which is written as
+$\hat{R}_{\text{clean}}(f)$ $=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}(f,\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]},\quad$
+$\text{where}\,\,\mathrm{Criterion}(\cdot)=\left\{\begin{array}[]{lr}1,\,\,\text{If selected},&\\ 0.&\end{array}\right.$ (4)
+In this unified formula, the classifier intended to be optimized is consistent with the classifier serving as the noise filter during the selection phase, which easily leads to the training bias.
+Empirical results in the section above show that directly optimizing ${\hat{R}}_{\text{clean}}(f)$ with an existing selection criterion is always biased, further resulting in inferior generalization performance.
+Overview.
+We propose to fulfill debiased learning by handling both the training and data bias.
+First, to solve the training bias, we design a novel network architecture called noIse-Tolreant Expert Model (ITEM).
 Compared with the widely applied double-branch network, our proposal exhibits greater robustness to noisy labels.
-Second, to solve the data bias, we propose a mixed sampling strategy that can mitigate the side-effect caused by the class-imbalanced set $\tilde{D}_{\text{clean}}$ while avoiding sparse representations.
+Second, to solve the data bias, we propose a weighted resampling strategy that can mitigate the side-effect caused by the class-imbalanced set $D_{\text{clean}}$ while avoiding sparse representations.
 3.1 ITEM: noIse-Tolerant Expert Model
-The training bias of sample selection stems from the way of a self-training manner.
+The training bias in sample selection arises from the self-training approach.
 Previous works [11, 21] always resort to the double-branch network to confront this problem.
 Concretely, the network is trained on the clean set that is selected by the other network.
-Motivated by Mixture-of-experts [41, 51], we propose a robust architecture that integrates multiple experts into the classifier, which independently conducts classifier learning and sample selection in a single network.
-Specifically, compared with a normal classification network with a classifier layer $f$, our network ITEM contains $f$ and a set of expert layers $\{g^{1},...,g^{m}\}$ with size $m$.
-For robust selection, we propose to conduct existing criteria on expert layers.
-Therefore, the training objective in our framework can be summarized as
-$\small\hat{R}_{\text{clean}}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\{g^{1},...,g^{m}\},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$
-Intuitively, the independent selection phase avoids the classifier layer selecting noisy samples and then updates its parameters on these data, which contributes to mitigating the training bias.
-We show the comparison of varying networks in Figure 3.
-Compared with the prevailing double-branch network, which is proved to be noise-robust in previous works [12, 10], our architecture contains two merits, including 1) resource-friendly training, in contrast to the feature extractor’s huge parameters, the last layer’s parameters for classification are significantly fewer.
-For example, the parameters of the classifier (or the expert) layer are less than 0.01 million, while the total parameters of a single ResNet-18 for CIFAR-10 are roughly 11.1 million.
+Motivated by Mixture-of-experts [31, 29], we propose a robust architecture that integrates multiple experts into the classifier, which independently conducts classification and sample selection.
+Specifically, compared with a normal classification network with a classifier layer $f$, our network ITEM contains $f$ and a set of additional expert layers $\{g^{1},...,g^{m}\}$ with size $m$.
+Each expert layer has the same dimension as the classifier layer, i.e., $\|f\|=\|g\|$.
+For robust selection, we propose to conduct existing criteria on expert layers instead of the classifier layer.
+Consequently, the training objective in our framework can be summarized as
+$\hat{R}_{\text{clean}}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}$ ($(\{g^{1},...,g^{m}\},\bm{x}_{i},{\tilde{y}}_{i})\cdot$
+$\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (5)
+Intuitively, the independent selection phase prevents the classifier layer from selecting noisy samples and updating its parameters based on them, significantly mitigating the training bias.
+In practice, a voting strategy is applied to ensemble the selection results of $m$ experts, where the sample selected by all experts is considered clean.
+It is noteworthy that the selection criterion $\mathrm{Criterion}(\cdot)$ is not restricted to a specific form (e.g., small-loss based or fluctuation-based), which highlights the applicability of this expert model as a backbone in sample selection.
+Advantages of the MoE structure.
+We compare different network structures in LNL in Figure 3.
+Compared with the prevailing double-branch networks, which have proved to be noise-robust in previous works [11, 21], our architecture offers two advantages, including 1) Greater robustness to noisy labels.
+Each expert exhibits different capabilities during training, resulting in a better inductive bias for both selection and prediction.
+Specifically, by ensembling the selections from different experts, the MoE minimizes negative interference from incompatible patterns (e.g., poor selection made by individual experts), thereby enhancing the framework’s robustness against noisy labels.
+2) Resource-friendly training.
+In contrast to the feature extractor’s huge parameters, the last layer’s parameters for classification are significantly fewer.
+The additional learnable parameters in our model are less than 2% of those in the double-branch network.
 Thus, our network is more friendly to GPU and computer memory.
-2) Greater performance, thanks to the property of fewer parameters, our model can integrate more experts, benefiting both selection and test accuracy by ensembling the results.
-Figure 4: The mapping function $\mathcal{S}^{\beta}(\cdot)$ with varying values of $\beta$. Note that $\mathcal{S}(\cdot)$ degrades into a transverse line when $\beta=1$.
-Remark.
-How to update parameters of the expert layer?
-All selection criteria depend on the model’s performance, including both small-loss family and fluctuation-based methods.
-Therefore, leveraging expert layers for selecting without updating parameters easily results in poor selection, even leading to a collapse.
-To fulfill accurate selection on expert layers, we propose a stochastic updating strategy, which randomly assigns a layer from the set of $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$ as the classifier layer $f^{c}$ and the rest are expert layers.
-Hence, Eq.
-(3.1) can be rewritten as
-$\small\hat{R}_{\text{clean}}(f^{c})=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\mathbbm{F}\backslash f^{c},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f^{c}(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (5)
-Since sufficient training processes, we consider that each layer in ITEM contains remarkable performance on selection and prediction.
-3.2 Mixed Sampling
-To overcome the data bias, we propose a Mixed Sampling strategy to simultaneously learn from head and tail classes in each training iteration.
-Specifically, supposing a clean set ${\tilde{D}}_{\text{clean}}$ is selected from the ${\tilde{D}_{N}}$, where ${\tilde{D}}_{\text{clean}}\subseteq{\tilde{D}_{N}}$.
-By calculating the number of instances from each class in ${\tilde{D}}_{\text{clean}}$ and conducting L1 normalization, a vector of class weights $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ can be obtained.
-The weight of class $k$ is written as
-$\bm{w}_{k}=\texttt{Normalize}(S_{k})=u_{k}\Big{/}\sum\nolimits_{i=1}^{K}u_{i},$ (6)
-where $u_{i}$ denotes the quantity of $i$-th class in ${\tilde{D}}_{\text{clean}}$.
-Based on the weight vector $\bm{v}$, we further propose a reversed weight vector that assigns a larger weight to tailed classes when sampling.
-Concretely, we construct a mapping function based on the probability density function (PDF) of a Beta distribution.
+3.2 Beta-Resampling
+Figure 4: Mapping function $\mathcal{S}^{\beta}(\cdot)$ with different values of $\beta$.
+To overcome the data bias, we propose a mixed beta-sampling strategy that promotes the sampling frequency for tail classes in the selected clean set when sampling.
+Thus, the model can learn from more knowledge from the tailed classes, contributing to mitigating the data bias.
+Specifically, suppose a clean set ${D}_{\text{clean}}$ is selected from the ${D_{N}}$, where ${D}_{\text{clean}}\subseteq{D_{N}}$.
+The set of samples labeled as class $k$ in ${D}_{\text{clean}}$ is ${D}_{\text{clean}}^{k}$.
+By introducing L1 normalization, a vector of class frequencies $\bm{v}$ can be obtained,
+$\bm{v}=[w_{1},w_{2},...w_{K}],\,\text{where}\,\,w_{k}=\frac{\|{D}_{\text{clean}}^{k}\|}{\|{D}_{\text{clean}}\|}.$ (6)
+Here, we can regard the class frequency as the sampling weight, i.e., the class with a larger frequency will be sampled with a greater weight during the data loading phase.
+To mitigate the data bias, which leads the model to learn from the majority classes mainly, we propose a mapping function that maps the lower frequency to a higher weight when data sampling.
+Concretely, this mapping function stems from the probability density function (PDF) of a Beta distribution.
 Normally, the Beta distribution’s PDF is controlled by two parameters $\alpha,\beta$.
-To satisfy both 1) the input interval of $\bm{w}_{k}$ belongs to $\{0,1\}$ and 2) the function contains the monotonically decreasing property, we thus fix $\alpha=1$ and adjust $\beta$ to achieve different slopes.
-Given a weight $\bm{w}$, the reversed weight ${\bm{\tilde{w}}}$ can be obtained by our mapping function, which is written as
-${\bm{\tilde{w}}}=\mathcal{S}^{\beta}(\bm{w})=\frac{1}{\text{B}(1,\beta)}\,(1-\bm{w})^{\beta-1},$ (7)
+To satisfy both 1) the input interval of ${w}_{k}$ belongs to $[0,1]$ and 2) the function contains the monotonically decreasing property, we thus fix $\alpha=1$ and adjust $\beta$ for different slopes.
+Given a weight $w_{i}$, the reversed weight ${\tilde{w}}_{i}$ is written as
+${\tilde{w}_{i}}=\mathcal{S}^{\beta}(w_{i})=\frac{1}{\text{B}(1,\beta)}\,(1-w_{i})^{\beta-1},$ (7)
 […]
-Algorithm 1 ITEM
-Input: Noisy training set $\tilde{D}_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyper-parameter $\beta$.
-1:  Initialize our network with one classifier layer and $m$ expert layers $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$.
-2:  for $t$ = 1, …, $T_{w}$ do
-3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
-4:  end for
-5:  for ${t}$ = 1, …, $T$ do
-6:     Select the clean set ${\tilde{D}}_{\mathrm{clean}}$ from $D_{N}$ based on $\textsf{Criterion}(\mathbbm{F})$.
-7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\mathrm{clean}}$.
-8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
-9:     while $i$ = 1, …, iterations do
-10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{v}$.
-11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{\tilde{v}}$.
-12:        Randomly sample $f^{c}$ from $\mathbbm{F}$.
-13:        Update the network parameters by minimizing $\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)}$.
-14:     end while
-15:  end for
-Training on the tail-focused training batch directly would result in the sparse representation of head classes [52].
-To avoid this problem, we separately conduct two times weighted sampling (according to the weighted and reversed weighted vector ${\bm{v,\tilde{v}}}$) in data loader phase.
-Then, we have two training mini-batches $B_{\bm{v}}$ and $B_{{\bm{\tilde{v}}}}$, which focus on head and tail classes, respectively.
-By leveraging MixUp strategy [22], the model can learn a more representative feature extractor on the mixed data $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$.
+The mapping function $\mathcal{S}^{\beta}(\cdot)$ for varying values of $\beta$ is illustrated in Figure 4.
+It can be observed that as the input weight increases, the output value decreases.
+Consequently, this mapping function enables the transformation of class frequencies in Eq.
+6 into a tail-focused weighted vector ${\bm{\tilde{v}}}=[\tilde{w}_{1},\ldots,\tilde{w}_{K}]$.
+By weighted sampling with ${\bm{\tilde{v}}}$, a tail-classes focused training batch $B_{{\bm{\tilde{v}}}}$ can be obtained.
+Training the model on $B_{{\bm{\tilde{v}}}}$ can effectively mitigate the data bias.
 Table 1: Test accuracy (mean±$\pm$±std) of methods using ResNet-18/34 on CIFAR-10/100.
-Note that $\dagger$,$\ddagger$ , ‡ and $\sharp$ denote three selection criteria, i.e., small-loss detection with loss threshold [10], small-loss detection with Gaussian Mixture Model [12], and fluctuation-based noise filtering [11].
+Note that $\dagger$,$\ddagger$ , ‡ and $\sharp$ denote three selection criteria, i.e., small-loss selection with loss threshold [11], small-loss selection with Gaussian Mixture Model [21], and fluctuation-based noise filtering [39].
 Bold values denote the best the second best performance.
 Methods Sym. 20% Sym. 40% Inst. 20% Inst. 40% Avg.
-CIFAR-10 Cross-Entropy [53] 85.00 $\pm$ 0.43% 79.59 $\pm$ 1.31% 85.92 $\pm$ 1.09% 79.91 $\pm$ 1.41% 82.61
-MentorNet [54] 80.49 $\pm$ 0.11% 77.48 $\pm$ 3.45% 79.12 $\pm$ 0.42% 70.27 $\pm$ 1.52% 76.84
+CIFAR-10 Cross-Entropy 85.00 $\pm$ 0.43% 79.59 $\pm$ 1.31% 85.92 $\pm$ 1.09% 79.91 $\pm$ 1.41% 82.61
 JoCoR [37] 88.69 $\pm$ 0.19% 85.44 $\pm$ 0.29% 87.31 $\pm$ 0.27% 82.49 $\pm$ 0.57% 85.98
@@
 $\sharp$ ITEM (Ours) 95.26 $\pm$ 0.23% 92.81 $\pm$ 0.20% 95.80 $\pm$ 0.18% 93.13 $\pm$ 0.29% 94.25
-CIFAR-100 Cross-Entropy [53] 57.59 $\pm$ 2.55% 45.74 $\pm$ 2.61% 59.85 $\pm$ 1.56% 43.74 $\pm$ 1.54% 51.73
-MentorNet [54] 52.11 $\pm$ 0.10% 35.12 $\pm$ 1.13% 51.73 $\pm$ 0.17% 40.90 $\pm$ 0.45% 44.96
+CIFAR-100 Cross-Entropy 57.59 $\pm$ 2.55% 45.74 $\pm$ 2.61% 59.85 $\pm$ 1.56% 43.74 $\pm$ 1.54% 51.73
 JoCoR [37] 64.17 $\pm$ 0.19% 55.97 $\pm$ 0.46% 61.98 $\pm$ 0.39% 50.59 $\pm$ 0.71% 58.17
@@
 $\sharp$ ITEM (Ours) 77.19 $\pm$ 0.13% 74.90 $\pm$ 0.24% 76.91 $\pm$ 0.23% 71.44 $\pm$ 0.29% 75.11
-3.3 Practical Training
-Despite the training objective of selection-based methods being summarized as a joint framework (in Eq.
-(5)), the training procedure is practically decoupled into two stages that are iteratively proceeding, 1) select clean samples via the preset criterion and 2) train the network on the selected set.
-Our training manner also obeys this framework.
-Specifically, supposing a selection criterion $\textsf{Criterion}(\cdot)$ such as small-loss [10] or fluctuation-based noise filtering [11] and a network with $m$ size expert layer.
-In each training epoch, the clean set ${\tilde{D}}_{\text{clean}}$ is selected.
-To calculate the number of each class in ${\tilde{D}}_{\text{clean}}$, we have a weighted vector $\bm{v}$.
-Then, the reversed vector $\bm{\tilde{v}}$ is obtained via the mapping function (in Eq.
-(7)).
-In each training iteration, we respectively sample a head-focused batch $B{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ and a tail-focused batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\text{clean}}$ according to ${\bm{\tilde{v}}}$, where $b$ denote the size of the mini-batch.
-By randomly selecting a layer from $\mathbbm{F}$ as the classifier layer $f^{c}$, the parameter’s update can be achieved by minimizing the following training loss
-$L^{\mathrm{train}}=\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)},$ (8)
-[…]
+3.3 Stochastic Classifier Training
+Since the noise filter and the optimizer are disentangled in Eq.
+(3.1), largely mitigating the training bias, the parameter update of expert layers is unreachable.
+Considering almost all existing selection criteria rely on the model’s performance, dynamically updating the parameters of expert layers is necessary.
+To fulfill this goal, we propose a stochastic training strategy, which randomly assigns a layer from the set of $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$ as the classifier layer $f^{c}$ and the rest are expert layers.
+Hence, Eq.
+(3.1) can be rewritten as
+$\hat{R}_{\text{clean}}(f^{c})=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}$ $(\mathbbm{F}\backslash f^{c},\bm{x}_{i},{\tilde{y}}_{i})\cdot$
+$\mathcal{L}(f^{c}(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (8)
+Since the aim-to-optimized layer $f^{c}$ is initialized from the set $\mathbbm{F}$ in each iteration, the performance of each layer can be ensured under sufficient training processes.
+Despite a highly integrated framework in Eq.
+(3.3) proposed, the training manner is practically decoupled into three stages that are proceeding iteratively, 1) select clean samples via the preset criterion based on the ensemble result of all experts, 2) randomly select an optimization layer, and 3) train this layer on the selected set.
+Considering training on the tail-focused training batch directly would result in the sparse representation of head classes [55], we separately conduct two times weighted samples (according to the weighted and reversed weighted vector ${\bm{v,\tilde{v}}}$) in each data-loading phase, which focus on head and tail classes, respectively.
+By leveraging the MixUp strategy [53], the model can learn a more representative feature extractor on the mixed data.
+Specifically, in each training iteration, we respectively sample a head-focused batch $B_{\bm{v}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ and a tail-focused batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\text{clean}}$ according to ${\bm{\tilde{v}}}$, where $b$ denotes the size of the mini-batch.
+By randomly selecting a layer from $\mathbbm{F}$ as the optimized-classifier layer $f^{c}$, the parameter’s update can be achieved by minimizing
+$L^{\mathrm{train}}=\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\mathrm{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\mathrm{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)},$ (9)
+where $\mathrm{MixUp}(,)$ is represented as $\mathrm{MixUp}(a,b)=\gamma\cdot a+(1-\gamma)\cdot b$ and $\gamma$ is a trade-off coefficient randomly sampled from a beta distribution $\texttt{beta}(\delta,\delta)$.
+The algorithm flowchart of ITEM is shown in Algorithm 1 (see Appdx).
 4 Experiments
 Datasets.
-We assess the performance of ITEM on two noise-synthetic datasets CIFAR-10 and CIFAR-100 [48], and two real-world noisy datasets CIFAR-10N [57] and Clothing-1M [58].
-To manually generate the noise, we follow the work [8] and construct symmetric and instance-dependent label noise on CIFAR-10 and CIFAR-10.
-Specifically, given a noise ratio $\rho$, symmetric label noise randomly translates the true label to a wrong label with probability $\rho$, while instance-dependent label noise assumes a corresponding relationship exists in input features [59].
-In experiments, we set the $\rho\in\{20\%,40\%\}$.
-For testing real-world performance, we employ CIFAR-10N, a human-annotated dataset with varying noise conditions, including worst and random 1&2&3.
-Besides, Clothing-1M [58], a large-scale dataset that is collected from the website and contains one million training data with 14 categories, is also utilized.
-Compared methods.
-On CIFAR-10 & 100, we compare ITEM with prevailing methods that can be roughly divided into three parts, including 1) sample selection methods (small-loss), MentorNet [54], JoCoR [8], and Me-Momentum [16], 2) sample selection methods (fluctuation), SFT [11] and Late Stopping [19], 3) others, Cross-Entropy [53], Joint Optim [55], and PES [56].
-Note that all reported results in this paper are collected from SFT and Late Stopping.
-On the real-world dataset CIFAR-10N, we compare ITEM with Co-teaching+ [14], Peer Loss [60], CAL [61], DivideMix [12], ELR+ [15], and CORES* [59].
-The results are collected from the literature [57].
-Experimental setup.
+We assess the performance of ITEM on two noise-synthetic datasets CIFAR-10 and CIFAR-100 [18], two human-annotated datasets CIFAR10N and CIFAR100N [38], and three real-world noisy benchmarks including Clothing-1M [46], Food-101N [20], and Webvision [23].
+Baselines.
+On CIFAR-10 and CIFAR-100, we compare ITEM with prevailing methods that can be roughly divided into three parts, including 1) small-loss based selection, JoCoR [37], and Me-Momentum [3], 2) fluctuation-based selection, SFT [39] and Late Stopping [51], 3) others, Cross-Entropy, Joint Optim [34], and PES [4].
+Reported results in this paper are collected from SFT and Late Stopping.
+More information is shown in Appdx.
+[…]
+Methods CIFAR10N CIFAR100N
+Worst R1 R2 R3
+Co-teaching+ [50] 83.26 89.70 89.47 89.54 60.37
+Peer Loss [27] 82.00 89.06 88.76 88.57 57.59
+CAL [57] 85.36 90.93 90.75 90.74 61.73
+Late Stoping [51] 85.27 - - - -
+$\ddagger$ ITEM 91.15 95.12 95.03 95.17 69.47
+DivideMix [21] 92.56 95.16 95.23 95.21 71.13
+ELR+ [26] 91.09 94.43 94.20 94.34 66.72
+CORES* [8] 91.66 94.45 94.88 94.74 61.15
+DPC [58] 93.82 95.97 95.92 95.90 71.42
+$\ddagger$ ITEM* 93.14 96.08 96.54 96.71 72.40
+Implementation details.
+Our code implements utilize Pytorch 1.9.0 and all experiments are run on a single RTX 4090 GPU.
 We keep the convention from [39, 51] and adopt ResNet-18 and ResNet-34 for CIFAR-10 and CIFAR-100, respectively.
-For all noisy conditions on CIFAR-10 & 100, we leverage an SGD optimizer with the momentum 0.9 and the weight-decay 1e-3 to train our network.
+For all noisy conditions on CIFAR-10 & 100, we leverage an SGD optimizer with the momentum $0.9$ and the weight-decay $1\times 10^{-3}$ to train our network.
 The total training epoch is set as 200.
 The initial learning rate is 0.02 and decayed with the factor 10 at the 100-th and 150-th epoch.
-For real-world noise, ResNet-34 and a pre-trained ResNet-50 are used for Clothing-1M and CIFAR-10N, respectively.
-More details can be found in the Table in Appendix C.
-Hyperparameter setup.
-Our framework contains two main hyperparameters, i.e., the number of expert $m$ in the network architecture, and the slope parameter $\beta$ in the mapping function.
-We keep $\beta=2$ for all experiments and set $m=2,4$ for CIFAR and Clothing-1M, respectively.
-4.1 Comparison Results
-Results on CIFAR-10 and CIFAR-100.
-Since our framework is not restricted by different selection criteria, we test our proposal on three types of selection criterion, including small-loss with loss threshold [10], small-loss with Gaussian Mixture Model [12] and fluctuation-based noise filtering [11].
-Each trial is repeated five times with different seeds.
-Mean values with standard deviation are recorded.
+For real-world noise, we adopt a ResNet-34 for CIFAR-N and a pre-trained ResNet-50 for Clothing-1M and Food101N, and Inception-ResNet-V2 [33] for WebVision.
+Other training set-ups like the number of training epochs, the learning rate and its adjustment, the optimizer, and so on keep the same as [25].
+Hyper-parameters.
+Our framework contains two main hyper-parameters, i.e., the number of expert $m$ in the network architecture, and the slope parameter $\beta$ in the mapping function.
+We keep $\beta=2$ for all experiments and set $m=4$ for CIFAR and $m=2$ for three open-world datasets.
+4.1 Results on Synthetic Noisy Datasets
+We conduct comparison experiments with two synthetic noise types on CIFAR-10 & 100, where we manually construct different noisy types on these two datasets.
+Since our framework is agnostic to selection criteria, we test our proposal on three types of selection criteria.
+Experimental results are shown in Table 1.
 […]
-[…]
-Methods CIFAR-10N
-Worst R. 1 R. 2 R. 3
-CE 77.69 85.02 86.46 85.16
-Co-teaching+ 83.26 89.70 89.47 89.54
-Peer Loss 82.00 89.06 88.76 88.57
-CAL 85.36 90.93 90.75 90.74
-Late Stoping 85.27 - - -
-$\ddagger$ ITEM 91.15 95.12 95.03 95.17
-DivideMix 92.56 95.16 95.23 95.21
-ELR+ 91.09 94.43 94.20 94.34
-CORES* 91.66 94.45 94.88 94.74
-$\ddagger$ ITEM* 93.14 96.08 96.54 96.71
-[…]
-Methods Test acc. (%) Methods Test acc. (%)
-CE 64.54 Co-teaching 69.21
-Forward 69.84 SFT+ 75.08
-JoCoR 70.30 DMI 72.46
-Joint Optim 72.23 PES (semi) 74.85
-DivideMix 74.76 ELR+ 74.81
-$\ddagger$ ITEM 74.69 $\ddagger$ ITEM* 75.02
-Results on CIFAR-10N.
-We evaluate the effectiveness of ITEM on a human-annotated noisy dataset CIFAR-10N which contains four noisy conditions, i.e., Worst and Random 1&2&3 [57].
-Besides, following previous selection-based approaches [12], we discard labels of unreliable samples and train the model with them by semi-supervised learning (SSL) for greater generalization, which is represented as ITEM*.
-Appendix A shows the detailed algorithm flowchart for ITEM*.
-The results can be found in Table 2.
+Table 3: Test accuracy (%) comparison of previous methods on three real-world noisy benchmarks. Top-1 and Top-5 test accuracy are reported for the dataset WebVision.
+Methods Food101N Clothing1M WebVision
+Top-1 Top-5
+Co-teaching [11] 83.73 67.94 63.58 85.20
+JoCoR [37] 84.04 69.06 63.33 85.06
+CDR [43] 86.36 66.59 62.84 84.11
+ELR+ [26] 85.77 74.81 77.78 91.68
+DivideMix [21] 86.73 74.76 77.32 91.64
+SFT+ [39] - 75.00 77.27 91.50
+CoDis* [45] 86.88 74.92 77.51 91.95
+SURE [25] 88.00 75.10 78.94 92.34
+$\ddagger$ ITEM* 88.14 75.08 80.20 93.07
+4.2 Results on Real-World Noisy Datasets
+We test ITEM’s performance on two human-made noisy datasets and three website noisy datasets.
+Considering that semi-supervised learning can utilize the sample discarded by the sample selection framework to regularize the model’s learning [21], we integrate ITEM with the Pseudo-labeling technique [19] to further improve the performance.
+1) Results on CIFAR10N & 100N.
+The results are shown in Table 2.
 First, our method achieves remarkable performance on both two settings (w/ or w/o SSL).
@@
 Compared with Late Stopping, a newly proposed method, ITEM achieves 5.88% improvements on worst-labels.
-When training with SSL, ITEM is also superior to the previous SOTA method DivideMix.
-It is worth noting that even training without SSL, ITEM surprisingly gains greater performance than those methods that leverage SSL approaches.
-Results on Clothing-1M.
-Going beyond the small-scale benchmark such as CIFAR-10 & 100, we also test the performance of ITEM and ITEM* on a large-scale dataset Clothing-1M, which contains roughly 38.49% noisy labels.
+When training with SSL, ITEM is superior to the previous SOTA method, DivideMix.
+Even training without SSL, ITEM surprisingly gains greater performance than those methods that leverage SSL approaches.
+2) Results on Food101N, Clothing1M, and WebVision.
 The results are shown in Table 3.
-We can observe that both ITEM and ITEM* outperform existing methods by a meaningful margin.
-For example, ITEM* achieves 0.17% improvements compared with the second-best method PES.
-The result verifies that our method is effective for enhancing model generalization in large-scale real-world scenarios.
-[…]
-$\xi$ 10
-$\rho$ 10% 20% 30% 40% 50%
-CIFAR-10 Co-teaching 80.30 78.54 68.71 57.10 46.77
-Sel-CL+ 86.47 85.11 84.41 80.35 77.27
-RCAL 88.09 86.46 84.58 83.43 80.80
-Ours 87.32 86.31 84.91 84.72 81.04
-CIFAR-100 Co-teaching 45.61 41.33 36.14 32.08 25.33
-Sel-CL+ 55.68 53.52 50.92 47.57 44.86
-RCAL 57.50 54.85 51.66 48.91 44.36
-Ours 56.10 55.93 53.07 48.29 45.17
-Table 5: Ablation study of each component in $\ddagger$ ITEM with varying noise conditions.
-Methods CIFAR-10 CIFAR-100 CIFAR-10N
-Sym 40% Inst. 40% Sym 40% Inst. 40% Worst
-w/o robust multi-expert network (training on normal ResNet) 90.75 91.44 73.01 70.04 89.72
-w/o mixed sampling (training on a random sampled batch) 88.77 89.91 71.09 68.54 87.66
-w/o mixup (directly training on $B{{\bm{v}}}$ and $B{{\bm{\tilde{v}}}}$) 90.19 90.35 72.50 69.11 89.10
-Ours 92.81 93.13 74.90 71.44 91.15
-(a) The number of expert layers $m$.
-(b) The slope parameter $\beta$.
-Figure 5: Sensitivity analyses for hyperparameters selection.
+On Food101N, our proposal achieved the best performance, outperforming SURE (the method with the second-best performance) [25] by 0.14%.
+On Clothing1M, the performance of our proposal is not the best but considerably competitive.
+Compared with the best performance, the 75.1% test accuracy, ITEM obtained a closed score, i.e., 75.08%.
+On WebVision, we keep the convention from the work [45] that tests the performance on the validation set of WebVision.
+ITEM significantly improved Top-1 accuracy by nearly 1.3%.
+The result verifies that our method effectively enhances model generalization in large-scale real-world scenarios.
+Figure 5: Visualization of debias learning in a class-level.
+We selected a ResNet-18 as the backbone and compared class-level prediction results of various methods on CIFAR-100 with four noise types.
+“CE (clean)" denotes training the model on the completely clean set (50k samples in total).
+The index of classes is sorted according to the class-level accuracy.
+Figure 6: Robustness comparison of different networks under two selection criteria. Our proposed MoE structure significantly mitigates the adverse effect of noisy labels.
 4.3 More Analyses
-[…]
-Ablation study.
-Our framework mainly contains two modules, the robust network architecture and the mixed data sampling strategy.
-We conduct ablation studies on CIFAR benchmarks to evaluate the effectiveness of each component (shown in Table 5).
-First, compared with typical ResNet architecture, our expert-based network indeed makes an obvious contribution to mitigating the noise.
-The average improvement among five noisy settings roughly reaches 2%.
-Second, the mixed sampling strategy also promotes the performance of our network.
-Note that compared with training on randomly sampled batches, training on two weighted sampled batches gains greater generalization.
-Thus, each module in our learning framework is effective.
-Hyperparameter selection.
-There are two main hyperparameters $m$ and $\beta$ in our framework.
-We conduct ablation studies to select optimal values for experiments, where $m\in\{2,3,4,5,6,7,8\}$ and $\beta\in\{1.3,1.8,2,3,5,10\}$.
-The result is shown in Figure 5.
-For the number of expert layers, we can see that the best test accuracy increases with a larger value of $m$.
-When $m=4$, the performance reaches the peak and slightly fluctuates as $m$ increases.
-Meanwhile, the gap between the peak value and the least value is around 0.6% on CIFAR-10 and 0.9% on CIFAR-100, respectively.
-For the slope parameter in the mapping function, we observe that the performance of our framework is not sensitive to varying values of $\beta$.
-Therefore, we select $m=4$ and $\beta=3$ for all experiments.
-Figure 6: Comparisons of class-level prediction results on CIFAR-100 with 40% symmetric noise. Note that “CE (clean)" denotes training the model on the completely clean set (50k samples).
-Class-balanced learning.
+Debias learning for LNL.
 We visualize the test accuracy in each class to demonstrate that ITEM achieves relatively balanced performance on varying categories while fulfilling greater generalization.
 We plot comparison results in Figure 5.
-First, compared with CE (clean), which is trained on 50k clean samples, ITEM achieves almost unbiased prediction results.
+First, ITEM achieves almost unbiased prediction results compared to CE (clean), which is trained on 50k clean samples.
 Second, compared with existing methods, ITEM takes tail classes better into account since two weighted samplers.
-Therefore, ITEM obtains greater performance in these categories.
-The result demonstrates our proposal has considerable potential in fulfilling class-balanced learning while tackling label noise.
+Therefore, ITEM performs better in these categories.
+Besides, under 20% symmetric label noise, ITEM achieved better performance than the model trained on an absolutely clean set (see the red line vs.
+the black line).
+These results demonstrate that our approach has significant potential to address the implicit bias inherent in sample selection frameworks for noisy label learning.
+Robustness of MoE structure.
+Compared with existing robust network, our proposal MoE exhibits better robustness.
+To clearly show this merit, we evaluate two selection criteria and plot the training curve under three networks in Figure 6.
+For the small-loss criterion, which easily incurs accumulation errors (see the green line (a)), the performance improvement with the MoE structure is significant.
+In contrast to the double-branch structure, which experiences slight performance degradation when the learning rate decreases at the 100th epoch (see the orange line in (b)), the learning curve with the MoE structure is more robust and stable.
+Second, in terms of fluctuation-based selection, the MoE demonstrates greater robustness, more stable training curves, and better generalization performance in noisy settings.
+In addition, its plug-in plug-out capability further verifies the versatility of the MoE structure.
+Table 4: Ablation studies of each component in ITEM with varying noise conditions.
+Methods CIFAR-10 CIFAR-100 CIFAR10N CIFAR100N
+Sym 40% Sym 40% Worst Fine
+w/o MoE Network 90.75 73.01 89.72 68.19
+w/o Mixed Sampling 88.77 71.09 87.66 68.24
+w/o MixUp 91.19 72.50 89.90 69.10
+Ours 92.81 74.90 91.15 69.72
+4.4 Ablation Studies
+Effect of each component.
+Our framework mainly contains two modules: the robust network architecture and the mixed data sampling strategy.
+We conduct ablation studies on CIFAR benchmarks to evaluate the effectiveness of each component from the following two perspectives.
+1) Quantitative analysis.
+We split the main component in ITEM into three parts: robust MoE, Mixed sampling, and the MixUp strategy.
+By solely removing them from ITEM, we reported the performance of the model in Table 4.
+Compared with typical ResNet architecture, our expert-based network indeed makes an obvious contribution to mitigating the noise.
+The average improvement among five settings roughly reaches 2%.
+Second, the mixed sampling strategy also promotes the performance of our network.
+Compared with training on randomly sampled batches, training on two weighted sampled batches gains greater generalization.
+2) Representation visualization.
+Considering that training with different sampling leads the model to different representation performance, we compare the learned representation space of the test set of CIFAR-10 and show the results in Figure LABEL:fig:TSNE.
+We can see that the learned representation of ten classes on the normal sampling $B$ is incompact, while the decision boundary is not well generalized.
+By contrast, if the proposed reversed tail-focused batch $B_{\tilde{v}}$ is adopted, each cluster is more compact where the distance between each cluster is larger, which verifies the effectiveness of our proposal sampling strategy in mitigating data bias underlying the selection phase.
+Eventually, by leveraging the Mixup operation, a feature extractor with better representation performance is obtained, as shown in the last plot.
+Figure 8: Sensitivity analyses of two hyper-parameters in ITEM. A ResNet-18 is trained on CIFAR-10 with 40% sym. noisy labels.
+Sensitivity analyses.
+ITEM has two main hyper-parameters, $m$ and $\beta$.
+We conduct ablation studies to select optimal values for experiments.
+The result is shown in Figure 8.
+First, the number of expert layers $m$ is the most important parameter in our framework, which decides the overall network structure and computation costs.
+From the right figure, we can observe that the performance of ITEM gradually increases with the value of $m$ increases.
+When arriving at a peak, i.e., 93.55% top-1 accuracy at $m=5$, the performance will not increase if the value of $m$ becomes larger.
+Therefore, we recommend a larger value of $m$ given an unknown dataset.
+Second, the value of $\beta$ has slightly influenced the performance of our method, where we also prefer a larger value of $\beta$.
+5 Related Work
+There are three main research lines in sample selection against the incorrect information from noisy labels.
+1) Robust selection criteria.
+Previous sample selection methods normally exploit the memorization effect of DNNs, i.e., DNNs first memorize training data with clean labels and then those with noisy labels [11, 43].
+The small-loss criterion [11] is a typical method stemming from the memorization effect, which splits the noisy training set via a loss threshold.
+The samples with small losses are regarded as clean samples.
+Some improved criteria [21, 1] based on the small-loss criterion are proposed in later works.
+In addition to the family of small loss criterion, some methods motivated by prediction fluctuation are designed.
+These methods [44, 56] observe that the model tends to give inconsistent prediction results for noisy samples and thus filter noisy labels by identifying high-frequency fluctuation samples.
+2) Robust network architecture.
+Easily works based on constructing the noise transition matrix always resort to an additional adaptation layer at the top of the softmax layer [14] or design a new dedicated architecture [10, 49].
+Recently, a family of double-branch networks [11, 50, 37, 21] has been proposed, which integrates two networks with the same architecture and selects clean samples for another network.
+Besides, better performance would be achieved when ensembling prediction results from two networks [11].
+3) Robust training procedures.
+Overall, current methods in sample selection always resort to a self-training manner, i.e., iteratively conducting clean label selection and model retraining.
+However, the incorrect selection would degrade the subsequent model learning.
+Recently, some robust training manners were proposed.
+Concretely, [56] designed a framework based on curriculum learning [5], which starts with learning from clean data and then gradually moves to learn noisy-labeled data with pseudo labels.
+Relations to us. The advantages of our proposal compared with previous methods can be divided into the following three parts, which are summarized in Table 5 in Appendix.
+  • Flexibility of the network structure.
+To our knowledge, we are the first to apply a mixture-of-experts (MoE) structure to LNL.
+Despite the core idea of MoE and double-branch structure being similar (i.e.,ensembling prediction results and the selection result over different views), MoE exhibits more flexibility.
+Expanding the number of experts is simple and only increases slight parameter counts (0.01 million per expert).
+By contrast, the double-branch structure is hardly extended to a three- or four-branch structure since the number of learnable parameters.
+  • Expansibility of the framework.
+Compared with previous selection-based methods, our proposal is agnostic to the selection criteria, which can be easily integrated with different selection criteria and improves their performance.
+  • Debias learning.
+Previous mainly focus on the accumulation error, a training bias.
+In this paper, we empirically show the existence of data bias (i.e., the imbalanced dataset selected by current selection criteria) and design an expert-based sampling strategy, which mitigates this bias by adjusting selection probabilities and increasing the visibility of underrepresented classes.
 6 Conclusion
 In this paper, we disclosed the data bias, an implicit bias underlying the sample selection strategy.
-To solve the training and data bias simultaneously, we proposed ITEM, which integrates a noise-robust multi-experts network and a mixed sampling strategy.
+To solve the training and data bias simultaneously, we proposed ITEM that introduces a noise-robust multi-experts network and a mixed sampling strategy.
 First, our structure integrates the classifier with multiple experts and leverages expert layers to conduct selection independently.
@@
 Second, our mixed sampling strategy yields class-aware weights and further conducts weighted sampling.
-By using the mixup of two weighted mini-batches, we mitigate the issue of sparse representation.
 The effectiveness of ITEM in real applications is verified on diverse noise types and both synthetic and real-world datasets.
-Appendix for Debiased Sample Selection for Combating Noisy Labels
-A ITEM with a Semi-Supervised Learning Framework
-Considering the large scale of unreliable samples discarded, especially in considerable noise ratios, leveraging these samples through semi-supervised learning is necessary.
-Therefore, we provide a novel version of ITEM with stronger performance.
+Appendix A Visualization of Data Bias
+In this section, we conducted extensive experiments that discovered the existence of the data bias, e.g., the imbalanced subset is selected by current selection-based frameworks.
+Despite Figure 1 in the section on Introduction, we show more results in the following figure (See Figure 9).
+We can observe that data bias widely exists in selection-based frameworks no matter the different noise types and different datasets.
+Refer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
+Figure 9: Extended results of Figure 1. (a) Small-loss with threshold [11], (b) Small-loss with GMM [21], and (c) Fluctuation-based noise filtering [39].
+Table 5: We compare ITEM with various selection-based methods over four perspectives.
+Methods Network architecture Selection-criterion specific Training debias Data debias
+Co-teaching [11], CoDis [45] Double-branch Small-loss criterion
+JoCoR [37], DivideMix [21], ProMix [36] Double-branch Small-loss criterion
+SFT [39], Late Stopping [51] Single-branch Fluctuation-based selection
+ITEM (ours) Mixture-of-Experts Not specific
+Appendix B Algorithm of ITEM
+We provided the pseudocode for ITEM in Algorithm 1.
+Algorithm 1 ITEM
+Input: Noisy training set ${D}_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, and the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyper-parameter $\beta$.
+1:  Initialize our network with one classifier layer and $m$ expert layers $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$.
+2:  for $t$ = 1, …, $T_{w}$ do
+3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
+4:  end for
+5:  for ${t}$ = 1, …, $T$ do
+6:     Select the clean set ${D}_{\mathrm{clean}}$ from $D_{N}$ based on $\textsf{Criterion}(\mathbbm{F})$.
+7:     Calculate weighted vector $\bm{v}=[{w}_{1},...,{w}_{K}]$ on ${D}_{\mathrm{clean}}$.
+8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[{\tilde{w}}_{1},...,{\tilde{w}}_{K}]$ with Equation (7).
+9:     while $i$ = 1, …, iterations do
+10:        Sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${D}_{\mathrm{clean}}$ according to $\bm{v}$.
+11:        Sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${D}_{\mathrm{clean}}$ according to $\bm{\tilde{v}}$.
+12:        Sample $f^{c}$ from $\mathbbm{F}$.
+13:        Update the network parameters by minimizing the loss in Equation (9).
+14:     end while
+15:  end for
+ITEM with semi-supervised learning.
+Considering the large scale of unreliable samples discarded, especially in considerable noise ratios, leveraging these samples through semi-supervised learning (SSL) to regularize the model is necessary as well as effective.
+Thanks to the flexible training framework, any SSL method can be integrated into our proposal ITEM.
+Here, we provide a novel version of ITEM with stronger performance.
 The algorithm is shown in Algorithm 2.
@@
 5:  for ${t}$ = 1, …, $T$ do
-6:     Split $D_{N}$ into the clean set ${\tilde{D}}_{\rm clean}$ and the unreliable set ${\tilde{D}}_{\rm noise}$ based on $\textsf{Criterion}(\mathbbm{F})$, and discard labels in ${\tilde{D}}_{\rm noise}$.
-7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\rm clean}$.
-8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
+6:     Split $D_{N}$ into the clean set ${D}_{\rm clean}$ and the unreliable set ${\tilde{D}}_{\rm noise}$ based on $\textsf{Criterion}(\mathbbm{F})$, and discard labels in ${\tilde{D}}_{\rm noise}$.
+7:     Calculate weighted vector $\bm{v}=[{w}_{1},...,{w}_{K}]$ on ${\tilde{D}}_{clean}$.
+8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[{\tilde{w}}_{1},...,{\tilde{w}}_{K}]$ with Eq. (7).
 9:     while $i$ = 1, …, iterations do
-10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{v}$.
-11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x}^{\prime}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\rm clean}$ according to $\bm{\tilde{v}}$.
-12:        Draw a mini-batch $B=\{(\bm{x}^{\prime\prime}_{i})\}_{i=1}^{2b}$ from ${\tilde{D}}_{\rm noise}$ without labels.
-13:        Generate pseudo-labels, then have $B=\{(\bm{x}^{\prime\prime}_{i},{\tilde{y}^{\prime\prime}}_{i})\}_{i=1}^{2b}$, where ${\tilde{y}^{\prime\prime}}_{i}=\mathbbm{E}_{f\sim\mathbbm{F}}[f(\bm{x}^{\prime\prime}_{i})]$.
+10:        Randomly sample a batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${D}_{\rm clean}$ according to $\bm{v}$.
+11:        Randomly sample a batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x}^{\prime}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${D}_{\rm clean}$ according to $\bm{\tilde{v}}$.
+12:        Draw a mini-batch ${\hat{B}}=\{(\bm{x}^{\prime\prime}_{i})\}_{i=1}^{2b}$ from ${D}_{\rm noise}$ without labels.
+13:        Generate pseudo-labels for each sample in ${\hat{B}}$, then have ${\hat{B}}=\{(\bm{x}^{\prime\prime}_{i},{\tilde{y}^{\prime\prime}}_{i})\}_{i=1}^{2b}$, where ${\tilde{y}^{\prime\prime}}_{i}=\mathbbm{E}_{f\sim\mathbbm{F}}[f(\bm{x}^{\prime\prime}_{i})]$.
 14:        Randomly sample $f^{c}$ from $\mathbbm{F}$.
-15:        Update network parameters via minimizing $\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(B_{{\bm{v}}}\text{ and }B_{{\bm{\tilde{v}}}},B)\big{)}$.
+15:        Update network parameters via minimizing $\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(B_{{\bm{v}}}\text{ and }B_{{\bm{\tilde{v}}}},{\hat{B}})\big{)}$.
 16:     end while
 17:  end for
-B More experimental results
-B.1 Results on CIFAR-100N
-We additionally verify the effectiveness of $\ddagger$ ITEM on another human-annotated dataset, CIFAR-100N, whose noise ratio is roughly 40.2%.
-We adopt the same settings on CIFAR-100 (see Table 8), and the results are shown in Table 6.
-Our proposal ITEM achieves state-of-the-art performance on both supervised and semi-supervised frameworks.
-w/o semi w/ semi
-CE 55.50 $\pm$ 0.66% ELR + 66.72 $\pm$ 0.07%
-Co-teaching 60.37 $\pm$ 0.27% CORES* 55.72 $\pm$ 0.42%
-SOP+ 67.81 $\pm$ 0.23% DivideMix 71.13 $\pm$ 0.48%
-ITEM 69.47 $\pm$ 0.18% ITEM* 72.40 $\pm$ 0.19%
-Table 6: Test accuracy (%) on the test set of CIFAR-100N.
-B.2 Visualization
-We verify the effectiveness of $\ddagger$ ITEM on varying noisy conditions and show the following noise-robust results.
-Figure 7: Visualization of class-level selection performance of ITEM on CIFAR-100 with 60% symmetric label noise.
-As the training process proceeds, ITEM gradually improves the F-score in all classes instead of only in head classes.
-(a) Training on randomly sampled mini-batch $B$, the best test accuracy is 88.42%.
-(b) Training on $B_{\tilde{v}}$ only, the best test accuracy is 90.17%.
-(c) Training on $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$, the best test accuracy is 93.04%.
-Figure 8: Visualizing feature representations on the CIFAR-10 test set by T-SNE. The noise setting is 40% symmetric noisy labels.
-(a) CIFAR-10 $w/$ Inst. 40%
-(b) CIFAR-100 $w/$ Inst. 40%
-Figure 9: Visualization of training curves under three settings.
-Suppose a small-loss criterion with the loss threshold, we compare our methods (Ours) with the proposed multi-expert network and single network on two noise conditions.
-  • robust selection processes.
-As shown in Figure 7, the total selection performance increases as the training processes, while our framework has not degraded the selection performance on tail classes.
-  • robust feature representations.
-We visualize the learned representation of the test set on CIFAR-10.
-We compare three training manners and show the results in Figure 8.
-Compared with training on randomly sampled mini-batch and $B_{\tilde{v}}$, our mixed sampling strategy yields a feature extractor with a strong representation.
-As shown in Figure 7(c), the distance among varying class clusters is closer, reflecting that the classifier learns more robust decision boundaries.
-  • robust training curves.
-We conduct ablation studies to verify the effectiveness of our proposed robust network architecture.
-The results on two noisy settings are shown in Figure 9.
-We can see that compared with a single network (a normal ResNet structure), our expert network achieves more robustness.
-Meanwhile, the mixed sampling strategy also contributes to improving generalization performance.
-Methods Stanford Dogs CUB-200-2011 Avg
-Sym 40% Asy 30% Sym 40% Asy 30%
-CE 51.42 58.08 64.01 56.02 57.38
-Label Smooth 70.22 64.99 54.39 56.80 61.60
-Conf. Penalty 68.69 64.50 52.40 54.33 59.98
-Co-teaching 49.15 50.50 46.57 50.60 49.21
-JoCoR 49.62 53.59 52.64 51.70 51.89
-CE + SNSCL 75.27 64.49 68.83 61.48 67.52
-$\ddagger$ ITEM 74.92 68.19 70.04 66.58 69.93
-Table 7: Test accuracy (%) on two fine-grained benchmarks with two noisy settings. The best and the second best performances are highlighted with Bold and underline, respectively.
-Table 8: Detailed settings about training procedure of ITEM.
-CIFAR-10 CIFAR-100 CIFAR-10N (all conditions) Clothing-1M
-Statistic data
- Class number 10 10 10 14
- Training size 50,000 50,000 50,000 1,000,000
- Testing size 10,000 10,000 10,000 10,000
+Appendix C Experimental Settings
+In this section, we reported the statistical data of seven datasets utilized in this paper and the whole experimental setup in Table 6.
+C.1 Dataset
+CIFAR-10 and CIFAR-100 [17] dataset consists of 60,000 images of 10 and 100 categories. Each image is with a resolution of 32 $\times$ 32.
+CIFAR10N and CIFAR100N [38] also consists of 60,000 images.
+Meanwhile, the noise labels in these two are human-made with varying noise conditions, including worst and random 1&2&3 for CIFAR10N and fine noise for CIFAR100N.
+Clothing1M [46] is a large-scale dataset that is collected from real-world online shopping websites.
+It contains 1 million images of 14 categories whose labels are generated based on tags extracted from the surrounding texts and keywords, causing huge label noise.
+The estimated percentage of corrupted labels is around 38.46%.
+A portion of clean data is also included in Clothing1M, which has been divided into the training set (903k images), validation set (14k images), and test set (10k images).
+We resize all images to $256\times 256$ as in [21] and then random crop to 224 $\times$ 224.
+[…]
+WebVision [23].
+Following [26], we use a mini version of WebVision where only the top 50 classes are utilized.
+This mini WebVision dataset contains approximately 66 thousand images and the corresponding noise ratio is roughly 20%.
+C.2 Other Baselines
+On two human-annonated datasets CIFAR10N and CIFAR100N, we compare ITEM with Co-teaching+ [50], Peer Loss [27], CAL [57], DivideMix [21], ELR+ [26], CORES* [8], and DPC [58].
+The results are collected from the literature [38].
+On three real-world noisy datasets, we compared ITEM with SURE [25], Co-teaching [11], MentorNet [12], JoCoR [37], CDR [43], ELR [26], DivideMix [21], SFT [39], CoDis [45], SURE [25].
+Table 6: Experimental settings about training procedure of ITEM.
+Datasets CIFAR-10 CIFAR-100 CIFAR-10N CIFAR-100N Clothing-1M Food-101N WebVision
+Class number 10 100 10 100 14 101 50
+Training size 50,000 50,000 50,000 50,000 1,000,000 310,009 66,000
+Testing size 10,000 10,000 10,000 10,000 10,000 25,250 2,500
 Training procedure
- Network ResNet-18 ResNet-34 Pre-trained ResNet-50
- Batch size 64 100
- Epoch 200 10
- Warmup epoch 10 30 10 1
+Network ResNet-18 ResNet-34 ResNet-50 InceptionResNetV2
+Batch size 64 100 32
+Epoch 200 10 100 30
+Warmup epoch 10 30 10 30 1 5 3
+Learning rate (LR) 0.02 0.001 0.02
+Weight decay 1e-3 5e-4
+LR scheduler divide 10 at [100,150]th epoch divide 10 at 5th epoch divide 10 at 50th epoch divide 10 at 20th epoch
 Optimizer SGD
 Momentum 0.9
- Learning rate (LR) 0.02 0.001
- Weight decay 1e-3 5e-4
- LR scheduler divided by 10 at [100,150]-th epoch divided by 10 at 5-th epoch
 Hyperparameters
- the num. of experts $m$ $m=4$ $m=2$
+experts number $m$ $m=4$ $m=2$ $m=3$
 the slope parameter $\beta$ $\beta=3$
-B.3 Results on Fine-Grained Noisy Settings
-The literature [5] pointed out that noisy label’ generation is strongly associated with fine-grained datasets.
-Following settings in SNSCL [5], we conduct experiments on four noisy fine-grained datasets, including Stanford Dogs [65] and CUB-200-2011 [66].
-A ResNet-18 pre-trained on ImageNet with three experts ($m=3$) is utilized.
+[…]
+Imbalanced ratio $\xi$ 10 100
+Noise ratio $\rho$ 10% 20% 30% 40% 50% 10% 20% 30% 40% 50%
+CIFAR-10 CE 80.41 75.61 71.94 70.13 63.25 64.41 62.17 52.94 48.11 38.71
+Co-teaching [11] 80.30 78.54 68.71 57.10 46.77 55.58 50.29 38.01 30.75 22.85
+RoLT [41] 85.68 85.43 83.50 80.92 78.96 73.02 71.20 66.53 57.86 48.98
+Sel-CL+ [22] 86.47 85.11 84.41 80.35 77.27 72.31 71.02 65.70 61.37 56.21
+CurveNet [13] 87.12 85.02 84.39 82.99 78.57 76.55 74.19 71.90 67.20 64.83
+RCAL [54] 88.09 86.46 84.58 83.43 80.80 78.60 75.81 72.76 69.78 65.05
+Ours 87.32 86.31 84.91 84.72 81.04 78.39 76.21 74.19 70.99 67.28
+CIFAR-100 CE 48.54 43.27 37.43 32.94 26.24 31.81 26.21 21.79 17.91 14.23
+Co-teaching [11] 45.61 41.33 36.14 32.08 25.33 30.55 25.67 22.01 16.20 13.45
+RoLT [41] 54.11 51.00 47.42 44.63 38.64 35.21 30.97 27.60 24.73 20.14
+Sel-CL+ [22] 55.68 53.52 50.92 47.57 44.86 37.45 36.79 35.09 31.96 28.59
+CurveNet [13] 55.96 54.60 51.28 48.10 44.61 40.91 39.71 35.90 32.09 29.71
+RCAL [54] 57.50 54.85 51.66 48.91 44.36 41.68 39.85 36.57 33.36 30.26
+Ours 56.10 55.93 53.07 48.29 45.17 44.00 41.90 40.44 37.28 34.80
+Appendix D More Experimental Results
+D.1 Results on Imbalanced Noisy Labels
+[…]
+Dataset construction.
+Let $T_{ij}(\bm{x})$ denote a probability that the true label $i$ of the instance $\bm{x}$ is corrupted to the noisy label.
+Therefore, given a noise ratio $rho$, the noise transition matrix satisfies that $T_{ij}(\bm{x})=\rho$ if $i=j$, otherwise $T_{ij}(\bm{x})=\frac{1}{C-1}\rho$, where $C$ denotes the number of classes.
+We first construct an imbalanced dataset on CIFAR with different imbalanced ratios $\xi=\{10,100\}$ and then corrupt the labels with varying noisy ratio $\rho=\{10\%,...,50\%\}$.
+We compare our proposal ITEM with various methods in LNL and imbalanced LNL, including Co-teaching [11], RoLT [41], Sel-CL+ [22], and RCAL [54].
+Comparison results are shown in Table 7.
+Weighted sampling, which mitigates data bias by adjusting selection probabilities and increasing the visibility of underrepresented classes during training, helps ITEM tackle long-tailed classification tasks.
+On both imbalanced noisy datasets, ITEM exhibits greater generalization under a high noise ratio.
+For example, ITEM outperforms RCAL by 1.29% and 0.24% under 40% and 50% noise ratio, respectively.
+On CIFAR-100 with an imbalanced ratio $\xi=100$, ITEM consistently achieves the best generalization performance over other comparison methods.
+Under an extreme noise rate $\rho=50\%$, ITEM achieves more than 4.5% improvement of top-1 test accuracy.
+Therefore, we believe that ITEM can also fulfill strong robustness even under imbalanced real-world noisy conditions.
+D.2 Visualization of Class-Level Selection
+As shown in Figure 10, we further visualize the class-level selection performance of our proposal.
+We can see that the total selection performance increases as the training progresses.
+Further, in the tailed classes where the value $F_{k}$ in other selection frameworks would decrease, ITEM gradually improves the F-score in all classes instead of only in head classes, demonstrating that ITEM effectively mitigates the data bias.
+Figure 10: Data debias on selection. We plot the class-level selection performance of ITEM on CIFAR-100 with 60% symmetric label noise.
+Table 8: Comparison of total training time in hours on CIFAR10N with Worst noise labels with varying experts number.
+Metric Single Network Double-branch Network Ours: MoE
+CE Co-teaching DivideMix ITEM ITEM ITEM* ITEM*
+[11] [21] ($m=3$) ($m=5$) ($m=3$) ($m=5$)
+Test accuracy (%) $\uparrow$ 77.69 83.26 92.56 91.15 91.40 93.14 93.32
+Times (hours) $\downarrow$ 1.6 4.2 4.4 1.6 1.7 1.8 1.8
+Table 9: Test accuracy (%) on two fine-grained benchmarks with two noisy settings. The best and the second best performances are highlighted with Bold and underline, respectively.
+Methods Stanford Dogs CUB-200-2011 Avg.
+Sym 40% Asy 30% Sym 40% Asy 30%
+Cross-Entropy 51.42 58.08 64.01 56.02 57.38
+Confidence Penalty [30] 68.69 64.50 52.40 54.33 59.98
+Label Smooth [28] 70.22 64.99 54.39 56.80 61.60
+Co-teaching [11] 49.15 50.50 46.57 50.60 49.21
+JoCoR [37] 49.62 53.59 52.64 51.70 51.89
+SNSCL [40] 75.27 64.49 68.83 61.48 67.52
+Ours 74.92 68.19 70.04 66.58 69.93
+D.3 Training Time Analysis
+In Table 8, we compare the training times of ITEM and (semi) ITEM* with three typical methods, using a single Nvidia 4090 GPU.
+We can observe that the training time of these methods based on double-brach networks (such as Co-teaching and DivideMix) is always twice as slow as training on a single network.
+However, in our proposed MoE structure, increasing the number of experts number will not significantly increase the computation cost.
+Besides, integrating ITEM with a semi-supervised learning framework can bring considerable performance improvement with fewer extra costs.
+D.4 Results on Fine-Grained Noisy Labels
+Recently, one literature [40] pointed out that noisy label generation is strongly associated with fine-grained datasets.
+Considering the property of wide existence and strong meaningfulness of noisy fine-grained classification, following settings in SNSCL [40], we conduct experiments on four noisy fine-grained datasets, including Stanford Dogs [15] and CUB-200-2011 [42].
+Given the selection criterion GMM-based small loss selection, we fine-tuned a (pre-trained) ResNet-18 with three experts.
 We synthetically construct two noise types, symmetric and asymmetric noise labels.
@@
 Our method achieves significant generalization performance even when tackling fine-grained noisy classification and consistently outperforms other methods by a large margin.
-We consider the MixUp operation to be the primary contribution to the advanced performance on fine-grained datasets since it encourages the model to learn from structured data instead of the unstructured noise [67].
-C Experimental Settings
-We reported the statistical data of four datasets and the whole experimental setup in Table 8. Besides, the data augmentation strategy for different datasets can be divided into two parts.
-  • For CIFAR-10, CIFAR-100, and CIFAR-10N, randomly cropping and horizontally flipping are adopted.
-  • For Clothing-1M, we first resize the image to 256x256 and then randomly crop it to 224x224. Besides, the horizontally flipping is adopted.
-Meanwhile, we detailedly exhibit the noise generation process by the noise transition matrix.
-As shown in Figure 10, given a noise ratio $\rho$, the noise transition matrix $T$ is represented as $T_{ij}(\bm{x})=P({\tilde{y}}=j|y=i)$, where $T_{ij}$ denotes that the label of the instance transits from a clean label $i$ to a noisy label $j$.
-The generation process of instance-dependent label noise can be found in the work [23, 11].
-Figure 10: Visualization of the transition matrix for symmetric (left) and asymmetric (right) noise on a ten classification task when the noise ratio is $40\%$.
-C.1 Settings for Learning with Noisy Labels on Imbalanced Datasets
-We obey the same manner of imbalanced noise construction from RCAL [62] and conduct experiments on synthetic CIFAR-10 and CIFAR-100.
-Concretely, we set the batch size as 64 and run 200 epochs for both CIFAR-10 and CIFAR-100.
-The learning rate is set as 0.02 and reduced by a factor of 10 at {100, 150}-th epoch.
-A ResNet-32 network with 4 experts is utilized.
-Besides, for a fair comparison, we adopt the same augmentation policy SimAug [68] as RCAL.
+For example, under Stanford Dogs with 30% asymmetric noise, ITEM improves the best method (label smooth) by more than 4%.
+We consider the MixUp operation to be the primary contribution to the advanced performance on fine-grained datasets for two reasons, 1) it encourages the model to learn from structured data instead of the unstructured noise [1], and 2) it encourages the model to learn a more generalized decision boundary, which largely benefits the fine-grained classification task.